### v1

# Free energy of the gas of spin 1/2 fermions beyond the second order and the Stoner phase transition

Abstract
In the previous work we have developed a systematic thermal (imaginary time) perturbative expansion and applying it to the relevant effective field theory computed, up to the second order in the interaction, the free energy $F$ of the diluted gas of (nonrelativistic) spin $1/2$ fermions interacting through a spin-independent repulsive two-body potential. Here we extend this computations to higher orders: assuming that the only relevant parameter specifying the interaction potential is the $s$-wave scattering length $a_{0}$, we compute the complete order $(k_{\rm F}a_{0})^{3}$ ($k_{\rm F}$ is the Fermi wave vector) contribution to the system’s free energy as a function of the numbers $N_{+}$ and $N_{-}$ of spin up and spin down fermions (i.e. as a function of its polarization) and the temperature $T$. We also extend the computation beyond a fixed order by resumming the contributions to $F$ of two infinite sets of Feynman diagrams: the so called particle-particle rings and the particle-hole rings. We find that including the second one of these two contributions has a dramatic consequence for the transition of the system from the paramagnetic to the ferromagnetic phase (the so called Stoner phase transition): in this approximation the phase transition simply disappears.
Keywords: Diluted gas of interacting fermions, effective field theory, itinerant ferromagnetism, phase transitions.
1. Introduction
In the recent years some progress has been achieved in the computation of equilibrium zero temperature properties of the gas of $N$ (nonrelativistic) fermions interacting with each other through a spin independent repulsive potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$. It was mainly related to the application to this classic [1, 2, 3, 4, 5] many-body quantum mechanics/statistical physics problem of general methods (see e.g. [6]) of the effective field theory. In this approach, initiated in the seminal paper [7] (see also [8]), the original spatially nonlocal potential $V_{\rm pot}$ assumed to be characterized by a length scale $R$ is replaced by the a priori infinite series of local interactions (written here using the standard second quantization formalism - see e.g. [9])
$\hat{V}_{\rm int}=C_{0}\!\int\!d^{3}{\mathbf{x}}~\!(\hat{\psi}_{+}^{\dagger}\hat{\psi}_{+})(\hat{\psi}_{-}^{\dagger}\hat{\psi}_{-})+\hat{V}_{\rm int}^{(C_{2})}+\hat{V}_{\rm int}^{(C_{2}^{\prime})}+\dots,$ (1)
of decreasing length dimension. The coefficients (couplings) $C_{0}$, $C_{2}$, $C_{2}^{\prime},\dots$ of the interaction (1) can be then directly determined in terms of the quantities - the scattering lengths $a_{0},a_{1},\dots$ and the effective radii $r_{0},\dots$ - parametrizing the general expansion (in powers of the relative momentum) of the amplitude of the elastic scattering of two particles. Trading the (bare) couplings of (1) for these measurable quantities which characterize the underlying potential $V_{\rm pot}$ of the binary interactions has also the effect of removing ultraviolet (UV) infinities engendered by the locality of (1). The simplifications brought in by this approach allowed first to easily reproduce [7] those terms of the perturbative expansion of the ground state energy $E$ of the system of spin $s$ fermions with equal densities of different spin projections which in the past were obtained by more traditional (and requiring considerably more work) methods of many-body quantum mechanics and to extend [10] this computation to the fourth order in the systematic, organized by the power counting rules [7], expansion in powers (in higher orders modified also by logarithms) of the dimensionless product $k_{\rm F}R$ of the system’s overall Fermi wave vector
$k_{\rm F}=(6\pi^{2}n/g_{s})^{1/3}~\!,$ (2)
(here $n=N/V$ is the overall density of the gas of spin $s$ fermions and $g_{s}=2s+1$) and the characteristic length $R$. The same approach allowed also to compute the ground state energy of the system of spin $1/2$ fermions for different densities of fermions with the up and down spin projections [11] recovering the old result of Kanno [12] (obtained analytically for the hard core interaction potential) and to easily extend it [13] to fermions having spin $s$ greater than 1/2 (and, therefore, more possible spin projections). Finally in the works [14, 15] and [16] this computation has been extended up to the third order in the systematic expansion in powers of $k_{\rm F}R$.
These results allowed to investigate more quantitatively the phase transition, called also the Stoner transition, to the ordered state in which the densities $N_{+}/V$ and $N_{-}/V$ of fermions of opposite spin projections in the case of $s=1/2$ are not equal and the system exhibits a nonzero polarization $P\equiv(N_{+}-N_{-})/N$ which, according to the standard qualitative argument based on the positivity of the repulsive interaction energy and the Pauli exclusion principle, for sufficiently strong repulsion and/or sufficiently high overall density, should be energetically (at zero temperature) favored over the state with equal densities. In this regime the system should, therefore, exhibit the property called itinerant ferromagnetism. It has been found that the clear first order character of this transition predicted by the second order of the perturbative expansion (in agreement with the general quantitative arguments given long time ago in [17]) - at sufficiently large values of the “gas parameter” $k_{\rm F}a_{0}$ two symmetric minima of $E(P)$ are formed away from $P=0$ separated from the one at $P=0$ by a finite barrier and at $(k_{\rm F}a_{0})_{\rm cr}=1.054$ they become the global minima ($|P_{\rm cr}|=0.58$) - gets appreciably weakened ($|P_{\rm cr}|$ shifts significantly towards $P=0$ and the barrier becomes much lower) if the complete order $(k_{\rm F}a_{0})^{3}$ contribution is taken into account [14, 15]. This seemed to support the results of the work [18] in which a certain class of contributions to $E$ (arising from the so-called particle-particle ring diagrams) has been resummed to all orders in $k_{\rm F}a_{0}$ finding that in this approximation the transition is continuous - the new minima start to continuously move away from $P=0$ as the gas parameter crosses some critical value $(k_{\rm F}a_{0})_{\rm cr}\approx 0.8$ ($0.858$ if only 1 hole-hole N1N-1 particle-particle parts of the $N$-th order particle-particle ring diagrams and $0.79$ if the complete $N$-th order particle-particle ring diagrams are resummed; see also [20] for a refinement of this approach). It should be added that the predictions of [18], which qualitatively are supported also by the results of a different approach [19], seem to agree quite well with the results obtained with the help of the Monte Carlo simulations [21].
However if all scattering lengths $a_{\ell}$ and effective radii $r_{\ell}$ are of the same order of magnitude, $\sim\!R$, the complete (according to the power counting rules [7] which apply, strictly speaking, only to this case) order $(k_{\rm F}R)^{3}$ contribution to the ground state energy depends also on $a_{1}$ and $r_{0}$ and it has been shown in [16] that in this case the character of the phase transition (at zero temperature) predicted by this approximation depends on the relative magnitudes and signs of the parameters $a_{1}$, $r_{1}$ and $a_{0}>0$ (that is, on the more detailed characteristics of the underlying potential $V_{\rm pot}$). The computation in which to all orders resummed are only (some) contributions depending on the powers of $k_{\rm F}a_{0}$ corresponds rather to the situation encountered in physics of dilute atomic gases in which the $s$-wave scattering length $a_{0}$ (it can be of either sign) is made positive and very large compared to the remaining parameters ($|a_{0}|\gg R\sim|a_{1}|,|r_{1}|,\dots$) by exploiting properties of the Feshbach resonances (see e.g. [22]). In this case however the underlying interaction of fermions (atoms) is attractive and the scattering length $a_{0}$ is positive in the regime in which bound states composed of two fermions of opposite spins can form. The true ground state of the system is then very different than the one of nointeracting atoms (they may not be adiabatically connected to one another in the thermodynamic limit implicit in the field theory approach) which is used in perturbative computations performed within the effective field theory approach. Although the formulae for the system’s energy density $E/V$ obtained perturbatively (or with the help of a resummation) by expanding around the ground state of noninteracting fermions can seem to imply the transition to the ordered state, they apply at best to a metastable (from the thermodynamic point of view) state of the system and in real experiments the transition (which in principle could occur in a metastable state [23]) to the ferromagnetic state has in fact not been observed [24, 25, 26] due to the too rapid formation (at very low temperatures at which these experiments were carried out) of atomic dimers (bosons). It can be also remarked that in the situation in which the underlying interaction is attractive (despite giving rise to a positive $s$-wave scattering length) the mentioned qualitative argument for the occurrence of the transition no longer applies and the expectation that the transition should occur is mainly based on the textbook mean field correction to the energy density [4, 5] (equivalent to the first order correction of the perturbative expansion) which depends only on the $s$-wave scattering length $a_{0}$.
Computation of the ground state energy density $E/V$ as a function of the densities of fermions with different spin projections allows only to investigate the equilibrium properties of the system of interacting fermions at zero (or very low) temperatures. It is however of interest to determine its behavior also at nonzero temperatures. (In the context of the physics of atomic gases it is physically clear that formation of atomic dimers, which at very low temperatures makes observation of the phase transition to the ordered phase impossible, at higher temperatures should be less important.) This requires computing one of the thermodynamic potentials of the system of interacting fermions. So far such a computation of the free energy $F$ has been done [27] only up to the second order, i.e. up to terms of order $(k_{\rm F}a_{0})^{2}$, using the old-fashioned thermal perturbation theory (see e.g. [28], par. 32) based on the ordinary second order Rayleigh - Schrödinger perturbative expression for energy levels entering the statistical sum. In [29] we have recovered this second order expression for $F$ using the systematic thermal perturbative expansion (exploiting the imaginary time formalism [9]) and reproduced the thermal characteristics of the Stoner phase transition it predicts (pointing out however problems - not discussed in [27] - with accurate numerical determination of the critical values of the polarization), but have encountered a technical problem which prevented us to immediately extend the computation to higher orders. Here we show how this problem can be resolved and applying the developed systematic thermal expansion to the first term of the effective field theory interaction (1) we derive the formulae allowing to compute numerically the complete order $(k_{\rm F}a_{0})^{3}$ corrections to the free energy $F(T,V,N_{+},N_{-})$. With additional work, including the contributions of the next two terms of (1) it would, of course, be possible to compute the complete order $(k_{\rm F}R)^{3}$ correction to the free energy.
Instead of completing the order $(k_{\rm F}R)^{3}$ corrections to the free energy which would allow to investigate in more details the thermal profile of the Stoner phase transition to the ordered state induced by truly repulsive spin independent two-body potentials $V_{\rm pot}$ which necessarily give rise to the parameters $a_{1}$ and $r_{0}$ of comparable magnitude to that of $a_{0}$ (all $\sim R)$, we in this paper profit from the possibility provided by the simpler structure of the terms of the expansion generated by the imaginary time formalism and resum to all orders in $k_{\rm F}a_{0}$ not only the contributions to the temperature-dependent free energy $F$ of the particle-particle ring diagrams (done for zero temperature in the papers [18, 20, 30]) but also of the particle-hole ring diagrams.
Including in the free energy $F(T,V,N,P)$ only the resummed contribution of the particle-particle ring diagrams we recover for $T=0$ the results of the works [18, 30] and can show how they are modified at nonzero temperatures not exceeding the Fermi temperature (we expect that the results obtained within the effective field theory should be valid for temperatures in this range). However as we show, inclusion of the contribution of the resummed particle-hole diagrams changes the situation drastically: the phase transition to the ordered state simply disappears (the minimum of $F$ is at $P=0$ for all values of the parameter $k_{\rm F}a_{0}$ and all temperatures). This is a somewhat surprising result and we comment on its possible meaning in the Conclusions.
2. The formalism
The natural statistical formalism in which to compute equilibrium properties of the gas of fermions the interaction of which preserve their spins and therefore the numbers $N_{\pm}$ of spin up and spin down particles, is the Grand Canonical Ensemble with two independent chemical potentials $\mu_{\pm}$. The relevant statistical operator is then
$\hat{\rho}={1\over\Xi_{\rm stat}}~\!e^{-\beta\hat{K}}~\!$, (3)
where $\beta\equiv 1/k_{\rm B}T$, with $T$ the temperature and $k_{\rm B}$ the Boltzmann constant, and
$\hat{K}=\hat{H}_{0}-\mu_{+}\hat{N}_{+}-\mu_{-}\hat{N}_{-}+\hat{V}_{\rm int}\equiv\hat{K}_{0}+\hat{V}_{\rm int}~\!.$ (4)
[…]
$\hat{K}_{0}=\sum_{\sigma=\pm}\!\int_{\mathbf{p}}(\varepsilon_{\mathbf{p}}-\mu_{\sigma})a^{\dagger}_{\sigma}({\mathbf{p}})a_{\sigma}({\mathbf{p}})~\!,$ (5)
with $\varepsilon_{\mathbf{p}}=\hbar^{2}{\mathbf{p}}^{2}/2m_{f}$. Standard systematic thermodynamical perturbative expansion [9, 29] gives the potential $\Omega$ in the form of the series in powers of the interaction $\hat{V}_{\rm int}$
$\Omega=\Omega^{(0)}-{1\over\beta}\sum_{N=1}^{\infty}{(-1)^{N}\over N!}\!\int_{0}^{\beta}\!d\tau_{N}\dots\!\int_{0}^{\beta}\!d\tau_{1}~\!{\rm Tr}\!\left(\hat{\rho}^{(0)}{\rm T}_{\tau}[\hat{V}_{\rm int}^{I}(\tau_{N})\dots\hat{V}_{\rm int}^{I}(\tau_{1})]\right)^{\rm con}.$ (6)
Here $\hat{V}_{\rm int}^{I}(\tau)=e^{\tau\hat{K}_{0}}\hat{V}_{\rm int}e^{-\tau\hat{K}_{0}}$ is the interaction operator in the (imaginary time) interaction picture, ${\rm T}_{\tau}$ is the chronological ordering and $\hat{\rho}^{(0)}$ is the statistical operator of the noninteracting system. The superscript “con” means that only connected contributions (Feynman diagrams) should be taken into account. The first term in (6) is the textbook [4, 5] grand thermodynamical potential of the nointeracting system
$\Omega^{(0)}(T,V,\mu_{+},\mu_{-})=-{V\over\beta}\sum_{\sigma=\pm}\int_{\mathbf{p}}\!\ln\!\left(1+e^{-\beta(\varepsilon_{\mathbf{p}}-\mu_{\sigma})}\right).$ (7)
Owing to the thermal analog of the Wick formula (see e.g. [9]) computation of the successive terms $\Omega^{(N)}$ of the expansion (6) reduces to drawing all possible connected Feynman diagrams with $N$ interaction vertices arising from $\hat{V}_{\rm int}$ joined by oriented lines and integrating over positions $\mathbf{x}$} and “times” $\tau$ labeling these vertices the corresponding products of free the thermal propagators
$-{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}(\tau_{2}-\tau_{1},{\mathbf{x}}_{2}-{\mathbf{x}}_{1})={1\over\beta}\sum_{n\in{\mathbb{Z}}}\int_{\mathbf{p}}e^{-i\omega_{n}^{\rm F}(\tau_{2}-\tau_{1})}~\!e^{i{\mathbf{p}}\cdot({\mathbf{x}}_{2}-{\mathbf{x}}_{1})}\left(-\tilde{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}(\omega_{n}^{\rm F},{\mathbf{p}})\right),$ (8)
the Fourier transform $\tilde{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}$ of which have the form [9]
$-\tilde{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}(\omega_{n}^{\rm F},{\mathbf{p}})={-\delta_{\sigma_{2}\sigma_{1}}\over i\omega_{n}^{\rm F}-(\varepsilon_{\mathbf{p}}-\mu_{\sigma})}~\!,$ (9)
associated with (oriented) lines connecting vertices of the diagram. The resulting “momentum” space Feynman rules are almost identical with the ordinary ones except that the integrations over frequencies (energies) are replaced by summations over the (fermionic) Matsubara frequencies $\omega_{n}^{\rm F}=(\pi/\beta)(2n+1)$, $n\in{\mathbb{Z}}$.
Applying this formalism with the interaction operator $\hat{V}_{\rm int}$ given by the first term of (1) one finds that the order $C_{0}$ term of the expansion (6) is simply given by
$\Omega^{(1)}=C_{0}V{\cal G}_{++}(0,{\mathbf{0}})~\!{\cal G}_{--}(0,{\mathbf{0}})~\!,$ (10)
with
${\cal G}_{\pm\pm}(0,{\mathbf{0}})=\int_{\mathbf{p}}\left[1+e^{\beta(\varepsilon_{\mathbf{p}}-\mu_{\pm})}\right]^{-1}~\!.$ (11)
Higher order contributions to the potential $\Omega$ can also be systematically computed. If $\hat{V}_{\rm int}$ in (4) were the true, spatially nonlocal, two-body interaction (corresponding to a two potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$, where ${\mathbf{x}}_{i}$ are the positions of fermions), the successive terms of the expansion (6) would be (ultraviolet) finite. If $\hat{V}_{\rm int}$ is the local interaction (1) of the effective theory, the successive terms of the expansion (6) involve ultraviolet divergences and have to be regularized. As in our previous works we will employ for this purpose the cutoff $\Lambda$ on the wave vectors of virtual particles. Finite (in the limit $\Lambda\rightarrow\infty$) contributions to the potential $\Omega$ (a physical quantity) are then obtained by systematically expressing the (bare) couplings of $\hat{V}_{\rm int}$ in terms of other measurable (physical) quantities. As it is customary, and in line (at least when the gas is very diluted) with the physical intuition that properties of the gas are mainly determined by elastic two-body collisions of its constituents, one expresses the couplings $C_{0}$, $C_{2}$, etc. of (1) in terms of the measurable quantities related to such a scattering process, namely in terms of the scattering lengths $a_{0}$, $a_{1}$, effective ranges $r_{0}$, etc. [7]. In this work we will only need to express the coupling $C_{0}$ in this way; the relevant formula obtained by matching the amplitude of the elastic fermion-fermion scattering computed perturbatively using the first term of the interaction (1) onto the general form of the same amplitude parameterized by $a_{0}$, $a_{1},\dots$ and $r_{0},\dots,$ reads [11, 14, 10]
$C_{0}={4\pi\hbar^{2}\over m_{f}}~\!a_{0}\left(1+{2\over\pi}~\!a_{0}\Lambda+{4\over\pi^{2}}~\!a^{2}_{0}\Lambda^{2}\dots\right)\equiv C_{0}^{\rm R}\left(1+{2\over\pi}~\!a_{0}\Lambda+{4\over\pi^{2}}~\!a^{2}_{0}\Lambda^{2}\dots\right).$ (12)
From the thermodynamic point of view much more convenient to work with than the potential $\Omega$ is the free energy $F$ which canonically depends on the variables $T$, $V$ and the particle numbers $N_{\pm}$ which are given by the derivatives
$N_{\pm}=-(\partial\Omega/\partial\mu_{\pm})_{T,V}~\!.$ (13)
In principle, in each order of the expansion (6) to construct the free energy one should invert the relations (13) to obtain the chemical potentials as functions of the particle numbers $N_{+}$ and $N_{-}$ and insert these in the formula $F=\Omega+\mu_{+}N_{+}+\mu_{-}N_{-}$. Thus the values of the chemical potentials $\mu_{+}$ and $\mu_{-}$ change with each successive order of the expansion and the procedure of constructing the free energy looks rather cumbersome. It turns out, however, that in the systematic expansion this procedure simplifies considerably: it amounts in effect to using the chemical potentials $\mu_{\pm}^{(0)}$ determined by inverting the formula (13) with $\Omega$ replaced by $\Omega^{(0)}$ given by (7) and omitting in the expansion (6) those diagrams which give vanishing contribution in computing the corrections $\Delta E$ to the ground state energy $E=E^{(0)}+\Delta E$ of the system of interacting particles using the ordinary Dyson expansion of the formula [7, 10, 11, 14, 15] ($T$ stands here for time, not for the temperature)
$\Delta E=\lim_{T\rightarrow\infty}{i\hbar\over T}~\!\langle 0|{\rm T}_{t}\exp\!\left({1\over i\hbar}\!\int_{-T/2}^{T/2}\!dt~\!V_{\rm int}^{I}(t)\right)\!|0\rangle~\!,$ (14)
in which $|0\rangle$ is the ground state of the noninteracting system of $N=N_{+}+N_{-}$ fermions. In the case of the interaction proportional to $C_{0}$ this has been explicitly demonstrated in [29] up to the third order of the perturbative expansion. This prescription is obviously consistent with the fact that in the zero temperature limit the corrections to the free energy obtained from the thermodynamic expansion should go over into the corresponding corrections to the ground state energy given by (14).
[…]
can be also easily taken into account in this formalism by including it in the free Hamiltonian $\hat{H}_{0}$ which amounts to shifting the chemical potentials $\mu_{\pm}\rightarrow\tilde{\mu}_{\pm}\equiv\mu_{\pm}\pm{\cal H}$ in $\hat{K}_{0}$ given by (4). The free energy is then given as the series
$F(T,V,{\cal H},N_{+},N_{-})=F^{(0)}+F^{(1)}+F^{(2)}+\dots,$ (16)
in which
$F^{(0)}(T,V,{\cal H},N_{+},N_{-})=\Omega^{(0)}(T,N,\tilde{\mu}_{+}^{(0)},\tilde{\mu}_{-}^{(0)})+(\tilde{\mu}_{+}^{(0)}-{\cal H})N_{+}+(\tilde{\mu}_{+}^{(0)}+{\cal H})N_{-}~\!,$ (17)
and $F^{(N)}(T,V,{\cal H},N_{+},N_{-})=\Omega^{(N)}(T,N,\tilde{\mu}_{+}^{(0)},\tilde{\mu}_{-}^{(0)})$ for $N=1,2,\dots,$ where, as explained above, in computing $\Omega^{(1)}$, $\Omega^{(2)}$ etc. one should take into account only those diagrams of the expansion (6) which give nonzero contributions to $\Delta E$. The (shifted) chemical potentials $\tilde{\mu}_{+}^{(0)}$, $\tilde{\mu}_{-}^{(0)}$ are given by
$\tilde{\nu}_{\pm}^{(0)}\equiv\tilde{\mu}_{\pm}^{(0)}/k_{\rm B}T=f^{-1}\!\left((1\pm P)\left({\varepsilon_{\rm F}(n)\over k_{\rm B}T}\right)^{3/2}\right),$ (18)
where $\varepsilon_{\rm F}(n)\equiv\hbar^{2}k_{\rm F}^{2}/2m_{f}$ and $f^{-1}(x)$ is the inverse of the monotonic function (mapping $\mathbb{R}$} onto $\mathbb{R}$$_{+}$) defined by the integral
$f(\nu)={3\over 2}\!\int_{0}^{\infty}\!d\xi~\!{\xi^{1/2}\over 1+e^{\xi-\nu}}~\!.$ (19)
If the computation of $F$ is restricted to the order $C^{\rm R}_{0}$ (i.e. to the order $k_{\rm F}a_{0}$) correction given, as follows from the formulated prescription and the result (10) and (11), by
$F^{(1)}=VC_{0}(N_{+}/V)(N_{-}/V)~\!,$ (20)
the condition for the minimum of $F$ with respect to $N_{+}$ and $N_{-}$ (at fixed $N_{+}+N_{-}=N$) which determines the system’s polarization $P$ takes the form ($t\equiv k_{\rm B}T/\varepsilon_{\rm F}$, $h\equiv{\cal H}/\varepsilon_{\rm F}$)
${8\over 3\pi}~\!(k_{\rm F}a_{0})~\!P+2h=t\left[f^{-1}\!\left({1+P\over t^{3/2}}\right)-f^{-1}\!\left({1-P\over t^{3/2}}\right)\right].$ (21)
If the asymptotic expansion
$f^{-1}(x)=x^{2/3}\left[1-(\pi^{2}/12)x^{-4/3}-(\pi^{4}/80)x^{-8/3}-(247\pi^{6}/25920)x^{-4}+\dots\right],$ (22)
valid for $x\gg 1$ (obtained by inverting the Sommerfeld expansion [28] of the function $f(\nu)$) is used, the formula (21) reproduces the textbook [4, 5] low temperature equilibrium condition (equivalent to the condition $\mu_{+}=\mu_{-}$) and leads to the well known prediction that the Stoner phase transition to the ordered state is continuous with divergent magnetic susceptibility characterized by the critical exponent $\gamma=1$ and a finite discontinuity of the heat capacity. (In fact, this continuous character of the transition is accidental: in the same approximation the transition is of first order in the system of spin $s>1/2$ fermions and/or if the space dimension is not 3.) If the correction $F^{(2)}$ is included, the transition becomes first order, at least at sufficiently low temperatures [27, 11, 13] in agreement with the arguments given in the past in [17]. However the computation of the complete order $(k_{\rm F}a_{0})^{3}$ correction to the ground state energy $E$ performed in [14, 15, 16] showed that the first order character of the transition at zero temperature, very clear in the second order approximation, is strongly weakened, at least as long as the contributions proportional to $k_{\rm F}r_{0}$ and $k_{\rm F}a_{1}$ (which, if the underlying interaction potential $V_{\rm pot}$ is “natural”, i.e. if all $a_{\ell}$, $r_{\ell}$, etc. are of the same order of magnitude, are of the same order, $(k_{\rm F}R)^{3}$, as the $(k_{\rm F}a_{0})^{3}$ correction) are not taken into account [16]. Below we extend the existing computations in two ways: we compute the complete (proportional to $(C_{0}^{\rm R})^{3}$, i.e. to $(k_{\rm F}a_{0})^{3}$) temperature dependent third order corrections to the free energy $F$ and, moreover, we show how to perform the ressumation of two infinite subsets of temperature dependent corrections to $F$ of which the first one is the finite temperature generalization of the subset of diagrams taken into account in (the last section of) ref. [18].
3. Order $(k_{\rm F}a_{0})^{2}$ and order $(k_{\rm F}a_{0})^{3}$ particle-particle corrections to $F$
We begin by recalling the computation of the order $(C_{0}^{\rm R})^{2}$ term $F^{(2)}$ performed in [29]. In agreement with the formulated prescription it is given by the single Feynman diagram shown in Figure 1. The corresponding analytical expression can be obtained by convoluting either two AA-”blocks” or two BB-”blocks” shown in the same Figure:
$F^{(2)}=-{1\over 2}~\!C_{0}^{2}V~\!{1\over\beta}\sum_{l\in\mathbb{Z}}\!\int_{\mathbf{q}}[A(\omega_{l}^{B},\mathbf{q})]^{2}=-{1\over 2}~\!C_{0}^{2}V~\!{1\over\beta}\sum_{l\in\mathbb{Z}}\!\int_{\mathbf{q}}[B(\omega_{l}^{B},\mathbf{q})]^{2}.$ (23)
To make the formulae resulting from convoluting $A$-blocks more transparent it will be convenient to introduce the following notation:
$N_{--}^{\mathbf{p}}\equiv n_{+}({\mathbf{p}})~\!n_{-}({\mathbf{q}}-{\mathbf{p}})~\!,$
$N_{++}^{\mathbf{p}}\equiv[1-n_{+}({\mathbf{p}})]~\![1-n_{-}({\mathbf{q}}-{\mathbf{p}})]~\!,$
$n_{\pm}({\mathbf{p}})=\left[1+\exp\{\beta(\varepsilon_{\mathbf{p}}-\tilde{\mu}_{\pm}^{(0)})\}\right]^{-1},$ (24)
$\{{\mathbf{p}}\}\equiv n_{+}({\mathbf{p}})+n_{-}({\mathbf{q}}-{\mathbf{p}})-1~\!,$
$[{\mathbf{p}}]\equiv\varepsilon_{\mathbf{p}}-\tilde{\mu}_{+}^{(0)}+\varepsilon_{{\mathbf{q}}-{\mathbf{p}}}-\tilde{\mu}_{-}^{(0)}~\!.$
At zero temperature $N_{--}^{\mathbf{p}}$ and $N_{++}^{\mathbf{p}}$ reduce respectively to $\theta(p_{{\rm F}+}-|{\mathbf{p}}|)\theta(p_{{\rm F}-}-|{\mathbf{q}}-{\mathbf{p}}|)$ and $\theta(|{\mathbf{p}}|-p_{{\rm F}+})\theta(|{\mathbf{q}}-{\mathbf{p}}|-p_{{\rm F}-})$, hence the subscripts -- and ++++. It is also easy to check that
$\{{\mathbf{p}}\}=N_{--}^{\mathbf{p}}-N_{++}^{\mathbf{p}}=N_{--}^{\mathbf{p}}\left(1-e^{\beta[{\mathbf{p}}]}\right).$ (25)
The form of the distribution functions $n_{\pm}({\mathbf{p}})$ plays the role only in the second one of these two identities.
                $A(\omega_{l+1}^{B},\mathbf{q})=$                      $^{q-k,~l-n}$                      $_{k,~n}$                              $B(\omega_{l}^{B},\mathbf{q})=$                                    $^{k,~n+l}$                                    $_{k+q,~n}$
Figure 1: The order $C_{0}^{2}$ diagram contributing to the thermodynamic potential $F$ of the gas of spin $1/2$ fermions and two “elementary” one-loop diagrams (AA- and BB-”blocks”) out of which the second order and those higher order (in the $C_{0}$ coupling) contributions which are taken into account in this work are composed. Solid and dashed lines denote propagators of fermions with the spin projections ++ and -, respectively.
In the introduced notation the $A$-block (obtained in [29]) takes the form
$A(\omega_{l}^{B},\mathbf{q})=\int_{\mathbf{p}}\!{\{{\mathbf{p}}\}\over i\omega_{l}^{B}-[{\mathbf{p}}]}~\!,$ (26)
After the sum in (23) over the bosonic Matsubara frequencies $\omega^{B}_{l}=(\pi/\beta)l$ is performed using the standard formulae [9, 29] one gets
${F^{(2)}\over V}=-{1\over 2}~\!C_{0}^{2}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!{\{{\mathbf{p}}_{1}\}\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\left({1\over 1-e^{\beta[{\mathbf{p}}_{1}]}}-{1\over 1-e^{\beta[{\mathbf{p}}_{2}]}}\right).$ (27)
Since the two terms are formally identical (after making in the integrals in one of the terms the interchange ${\mathbf{p}}_{1}\leftrightarrow{\mathbf{p}}_{2}$), one arrives, using (25), at the final form of $F^{(2)}/V$:
${F^{(2)}\over V}=C_{0}^{2}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!N_{--}^{{\mathbf{p}}_{1}}~\!{1\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}-C_{0}^{2}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!,$ (28)
in which $\{{\mathbf{p}}\}^{\rm sub}\equiv\{{\mathbf{p}}\}+1=n_{+}({\mathbf{p}})+n_{-}({\mathbf{q}}-{\mathbf{p}})$ and the integrals should be understood in the Principal Value sense. Notice that the denominators ${\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]$ do not depend on the chemical potentials. This profiting from the symmetry of the two terms of (27), seemingly not problematic, has indeed no consequences here but, as will be shown, in higher orders if applied blindly would lead to incorrect results.
The first term in (28) is divergent. The change of the variables ${\mathbf{q}}=2{\mathbf{s}}$, ${\mathbf{p}}_{1}={\mathbf{s}}-{\mathbf{t}}_{1}$, ${\mathbf{p}}_{2}={\mathbf{s}}-{\mathbf{t}}_{2}$ (the Jacobian equals 8) makes the innermost integral elementary and allows to write it in the form
${16\pi^{2}\hbar^{4}\over m^{2}_{f}}~\!a^{2}_{0}\left(1+{4\over\pi}~\!a_{0}\Lambda+\dots\right)\!\int_{\mathbf{s}}\!\int_{{\mathbf{t}}_{1}}\!8~\!n_{+}({\mathbf{s}}-{\mathbf{t}}_{1})~\!n_{-}({\mathbf{s}}+{\mathbf{t}}_{1})~\!{m_{f}\over 2\pi^{2}\hbar^{2}}\left(-\Lambda+{{\mathbf{t}}^{2}_{1}\over\Lambda}+\dots\right),$
after using (12).
Expressing $C_{0}$ similarly in the second term of the formula (28) and in $F^{(1)}$ given by (20), one finds that the divergent terms of order $a_{0}^{2}\Lambda$ cancel out and
${F^{(1)}+F^{(2)}\over V}=C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})-(C_{0}^{\rm R})^{2}\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}$
$-{16\hbar^{2}\over\pi m_{f}}~\!a_{0}^{3}\!\int_{\mathbf{s}}\!\int_{{\mathbf{t}}_{1}}8~\!n_{+}({\mathbf{s}}-{\mathbf{t}}_{1})~\!n_{-}({\mathbf{s}}+{\mathbf{t}}_{1})~\!(\Lambda^{2}-2{\mathbf{t}}_{1}^{2})\phantom{aaaaaaaaa}$~\! (29)
$-{64\pi\hbar^{4}\over m_{f}^{2}}~\!a_{0}^{3}\Lambda\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\ N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}+{\cal O}(1/\Lambda)~\!.\phantom{aaaaaaaa}$~\!
The first two terms constitute the complete, finite contribution to $F/V$ up to the order $(C_{0}^{\rm R})^{2}$; the remaining terms are formally of higher order and can be considered only after including other third and higher order contributions.
In [29] it has been found that it is convenient to evaluate (the finite part of) $F^{(2)}/V$ by substituting ${\mathbf{p}}_{1}={\mathbf{k}}_{1}$, ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$, ${\mathbf{p}}_{2}={\mathbf{p}}$ (the Jacobian is 1), replacing (by another change of the integration variable) $n_{-}({\mathbf{k}}_{1}+{\mathbf{k}}_{2}-{\mathbf{p}})$ with $n_{-}({\mathbf{p}})$ and then performing explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$. This allows to represent the order $(k_{\rm F}a_{0})^{2}$ contribution to $F$ in the form
${F^{(2)}\over V}=C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})~\!L({\mathbf{k}}_{1},{\mathbf{k}}_{2})~\!.$ (30)
The (dimensionless function) $L({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by the single integral
$L({\mathbf{k}}_{1},{\mathbf{k}}_{2})=-{C_{0}^{\rm R}m_{f}\over(2\pi)^{2}\hbar^{2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}\!\int_{0}^{\infty}\!dp~\!p~\![n_{+}(p)+n_{-}(p)]\ln\!\left|{(p-\Delta_{+})(p-\Delta_{-})\over(p+\Delta_{+})(p+\Delta_{-})}\right|,$ (31)
in which
$\Delta_{\pm}={1\over 2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|\pm{1\over 2}|{\mathbf{k}}_{1}-{\mathbf{k}}_{2}|~\!.$ (32)
In [29] we have checked that in the zero temperature limit, in which the Fermi distribution functions $n_{+}({\mathbf{p}})$ and $n_{-}({\mathbf{p}})$ are replaced by the step functions $\theta(p_{{\rm F}+}-|{\mathbf{p}}|)$ and $\theta(p_{{\rm F}-}-|{\mathbf{p}}|)$, this formula reproduces numerically the second order correction to the ground state energy computed first (analytically) by Kanno [12] and then recovered (semi-analytically) in several works (e.g. in [11, 13]) for all values of the polarization $P$. We have also analyzed the free energy $F$ with the corrections $F^{(1)}$ and $F^{(2)}$ included and recovered, up to uncertainties following from the finite precision of the (rather complicated) numerical evaluation of the relevant multiple integrals the characteristics of the phase transition to the ordered state (for temperatures $T<\varepsilon_{\rm F}/k_{\rm B}$) first obtained in [27].
Figure 2: The particle-particle and the particle-hole diagrams contributing in the order $C^{3}_{0}$ to the thermodynamic potential $F$.
If only the interaction term proportional to $C_{0}$ in (1) is taken into account, there are two Feynman diagrams contributing to the free energy $F$ in the third order. The first one, shown in the left panel of Figure 2, is termed the particle-particle ring diagram. Its contribution $F^{(3)pp}$ is given by the convolution of three $A$-blocks
${F^{(3)pp}\over V}={1\over 3}~\!C_{0}^{3}~\!{1\over\beta}\sum_{l}\int_{\mathbf{q}}\![A(\omega_{l}^{\rm B},\mathbf{q})]^{3}~\!.$ (33)
After decomposing the product of three $A$-blocks into simple fractions, performing the summation over the bosonic Matsubara frequencies $\omega_{l}^{\rm B}$ and then using the identities (25) one arrives at
${F^{(3)pp}\over V}={1\over 3}~\!C_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\int_{{\mathbf{p}}_{3}}\!\left(N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{\{{\mathbf{p}}_{3}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms}\right),$ (34)
where “two other terms” means the terms in which the role of ${\mathbf{p}}_{1}$ is played by ${\mathbf{p}}_{2}$ and ${\mathbf{p}}_{3}$. It is good to make at this point a contact with the contribution of this third order particle-particle ring diagram to the ground state energy density $E/V$ obtained in [14, 15] (and in [16]) to which the expression (34) should reduce in the zero temperature limit, that is when the Fermi distribution functions are replaced by the theta functions. The contribution to $E/V$ of the particle-particle diagram was in [14, 15] given by two terms (Eq. (21) in [15]) whereas here it is given by the single term (three seemingly identical terms). The equivalence of the two approaches is ensured by the algebraic, i.e. independent of the precise forms of $N_{--}^{\mathbf{p}}$ and $\mathbf{p}$}] (recall that $\{{\mathbf{p}}\}=N_{--}^{\mathbf{p}}-N_{++}^{\mathbf{p}}$), identity which results from the symmetrization:
$N_{--}^{{\mathbf{p}}_{1}}~\!{N_{--}^{\mathbf{p}_{2}}-N_{++}^{\mathbf{p}_{2}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{2}]}~\!{N_{--}^{\mathbf{p}_{3}}-N_{++}^{\mathbf{p}_{3}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{3}]}+{\rm two~other~terms}\phantom{aaaaaaaaaaaaaaaaaaaaaaaaaaaaa}$ (35)
$=\left(N_{--}^{\mathbf{p}_{1}}~\!{N_{++}^{\mathbf{p}_{2}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{2}]}~\!{N_{++}^{\mathbf{p}_{3}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{3}]}+N_{++}^{\mathbf{p}_{1}}~\!{N_{--}^{\mathbf{p}_{2}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{2}]}~\!{N_{--}^{\mathbf{p}_{3}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{3}]}\right)+{\rm two~other~terms}.$
After using the symmetry, i.e. taking only the content of the bracket and multiplying it by 3, it allows to rewrite the expression for $F^{(3)pp}/V$ as the sum of two terms which in the $T=0$ limit precisely reduce to the two terms, $G_{1}$ and $G_{2}$, which in [14, 15] contributed to $E/V$.
Naively, as all the three terms of (34) seem also identical, one is tempted to compute only one of them and multiply the result by three. $F^{(3)pp}/V$ would be in this way given by a single four-fold integral. This, as we have found, would lead to an incorrect result which in the zero temperature limit would not agree with the contribution of the particle-particle diagram to $E/V$ (this is precisely the problem which did no allow us to immediately extend the computation reported in [29]).
To understand the problem it is instructive to consider the triple integral
$\int_{0}^{1}\!dx\!\int_{0}^{1}\!dy\!\int_{0}^{1}\!dz\left({1\over(x-y)(x-z)}+{1\over(y-x)(y-z)}+{1\over(z-x)(z-y)}\right).$ (36)
The integrand is algebraically zero and the result of the integration should be zero too. Yet the integrand has (spurious) singularities and the integrals in (36), similarly as the ones encountered in the computation of $F^{(3)}$, should be understood in the Principal Value sense. If one naively says that the integrals of the three terms are equal and evaluates only one of them (multiplying it by 3) one will get
$3\int_{0}^{1}\!dx\!\left({\rm P}\!\int_{0}^{1}\!dy~\!{1\over x-y}\right)^{2}=3\int_{0}^{1}\!dx~\!\ln^{2}{1-x\over x}=3~\!{\pi^{2}\over 3}.$
The correct result (zero) is obtained if one first regularizes the integrand of (36) by setting $x\rightarrow x+i\epsilon$, $y\rightarrow y+2i\epsilon$, $z\rightarrow z+3i\epsilon$ (the sign of $\epsilon$ is irrelevant; the integrand is still algebraically zero but its singularities are now off the integration axes). It is then straightforward to find that the application of the Sochocki formula $1/(x\pm i0)=P(1/x)\mp i\pi\delta(x)$ to the regularized integral (36) leads to (the terms linear in the Dirac deltas neatly cancel out)
$3\int_{0}^{1}\!dx\!\left({\rm P}\!\int_{0}^{1}\!dy~\!{1\over x-y}\right)^{2}+\int_{0}^{1}\!dx\!\left(i\pi\!\int_{0}^{1}\!dy~\!\delta(x-y)\right)^{2}=0~\!.$
If the same procedure is applied to (34) one gets
${F^{(3)pp}\over V}=C_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!~N_{--}^{{\mathbf{p}}_{1}}\!\left[\left({\rm P}\!\int_{{\mathbf{p}}_{2}}~\!{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\right)^{2}+{1\over 3}\left(i\pi\!\int_{{\mathbf{p}}_{2}}\!\{{\mathbf{p}}_{2}\}~\!\delta([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}])\right)^{2}\right].$ (37)
One can now check that in the sum $F^{(1)}+F^{(2)}+F^{(3)pp}$ all the divergences (up to the order $(k_{\rm F}a_{0})^{3}$) cancel out (as will be seen, the contribution $F^{(3)ph}$ of the other diagram of Figure 2, which completes the order $(k_{\rm F}a_{0})^{3}$ contribution to $F^{(3)}$, is finite; this also follows from the computations of the order $(k_{\rm F}a_{0})^{3}$ corrections to $E/V$ performed in [14, 15]). Writing $\{{\mathbf{p}}_{i}\}=-1+\{{\mathbf{p}}_{i}\}^{\rm sub}$ in the first term in the square bracket in (37) allows to single out the divergent part of $F^{(3)}/V$. It is given by (to this order one can set in (37) $C_{0}=C_{0}^{\rm R}$; we also suppress the symbol $P$ of the Principal Value)
${F^{(3)pp}_{\rm div}\over V}={64\pi^{3}\hbar^{6}\over m_{f}^{3}}~\!a_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}N_{--}^{{\mathbf{p}}_{1}}\!\left(\int_{{\mathbf{p}}_{2}}~\!{1\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\right)^{2}\phantom{aaaaaaaaaaa}$
$-2~\!{64\pi^{3}\hbar^{6}\over m_{f}^{3}}~\!a_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}N_{--}^{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\int_{{\mathbf{p}}_{3}}~\!{1\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}~\!.$ (38)
Making now in the first term the change of the variables ${\mathbf{q}}=2{\mathbf{s}}$, ${\mathbf{p}}_{1}={\mathbf{s}}-{\mathbf{t}}_{1}$, ${\mathbf{p}}_{2}={\mathbf{s}}-{\mathbf{t}}_{2}$ (the Jacobian is 8) and performing the innermost integral (over $d^{3}{\mathbf{t}}_{2}$) one finds that it precisely cancels the entire middle line of (29). Moreover, after making similar changes of the variables in the last line of (29) and in the last term of $F^{(3)pp}_{\rm div}/V$ they too mutually cancel out. The remaining contribution of the left diagram of Figure 2 is, therefore, given by (37) with $\{{\mathbf{p}}_{2}\}$ in the first term (but not in the second one!) replaced by $\{{\mathbf{p}}_{2}\}^{\rm sub}$ and $C_{0}$ replaced by $C_{0}^{\rm R}$. Making as previously the change ${\mathbf{p}}_{1}={\mathbf{k}}_{1}$, ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$, ${\mathbf{p}}_{2}={\mathbf{p}}$ of the integration variables one can represent the contribution of the particle-particle order $(k_{\rm F}a_{0})^{3}$ diagram to $F/V$ in the form
${F^{(3)pp}_{\rm fin}\over V}=C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{1}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})\left[L^{2}({\mathbf{k}}_{1},{\mathbf{k}}_{2})+{1\over 3}\left(iL_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})\right)^{2}\right],$ (39)
where the function $L({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by (31) while the dimensionless function $L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by the finite integral
$L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})=\pi~\!{C_{0}^{\rm R}m_{f}\over(2\pi)^{2}\hbar^{2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}\!\int_{p_{\rm min}}^{p_{\rm max}}\!dp~\!p~\![n_{+}(p)+n_{-}(p)-1]~\!,$ (40)
in which $p_{\rm min}=|\Delta_{-}|$ and $p_{\rm max}=\Delta_{+}$ are determined by the condition that the zero of the argument of the Dirac delta treated as a function of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lies between 1-1 and +1+1. Since $\varepsilon_{\mathbf{p}}$ depends on $p^{2}$, the function $L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can, of course, be obtained in a closed analytic form. The expression (39) is finite (as indicates its superscript) the ultraviolet convergence [^3] The singularity introduced by the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|^{2}$ is superficial because for ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}={\mathbf{0}}$ vanishes also the logarithm in (31) while in (40) $p_{\rm min}=p_{\rm max}$. of the integrations being secured by the exponential suppression provided by the Fermi distribution functions $n_{\pm}$. We have also checked that the expression (39) evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-particle diagram of Figure 2 obtained in [14, 15].
4. Resummation of the contributions of the particle-particle diagrams
It turns out that the contribution to the free energy $F$ of the infinite series of Feynman diagrams composed of $N$-fold rings of the particle-particle $A$-blocks of Figure 1 can be summed in a closed form. Consider first the order $(k_{\rm F}a_{0})^{N}$ term of this series (the factor $(-1)^{N+1}$ is the same as in (6) - there are as many rearrangements of the $\hat{\psi}_{+}$ operators as of the $\hat{\psi}_{-}$ ones; the factor $1/N!$ in (6) is reduced to $1/N$ as there are (N1)!(N-1)! identical diagrams)
${F^{(N)pp}\over V}=(-1)^{N+1}~\!{C_{0}^{N}\over N}~\!{1\over\beta}\sum_{l}\int_{\mathbf{q}}[A(\omega_{l}^{\rm B},{\mathbf{q}})]^{N}~\!.$ (41)
Decomposing the product of the integrands of the NN AA-blocks (26) using the identity
$\prod_{i=1}^{N}{1\over x-a_{i}}=\sum_{n=1}^{N}\left(\prod_{j\neq n}^{N}{1\over a_{n}-a_{j}}\right){1\over x-a_{n}}~\!,$ (42)
and performing then the summation over the Matsubara frequencies one gets the integrand of the $(N+1)$-fold integral in the form
$\{{\mathbf{p}}_{1}\}\dots\{{\mathbf{p}}_{N}\}\sum_{n=1}^{N}\left(\prod_{j\neq n}^{N}{1\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]}\right){1\over 1-e^{\beta[{\mathbf{p}}_{n}]}}~\!,$ (43)
and finally, after using the relations (25), $F^{(N)pp}/V$ takes the form
${F^{(N)pp}\over V}=(-1)^{N+1}~\!{C_{0}^{N}\over N}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\dots\!\int_{{\mathbf{p}}_{N}}\sum_{n=1}^{N}N_{--}^{{\mathbf{p}}_{n}}\left(\prod_{j\neq n}^{N}{\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]+i(n-j)\epsilon}\right),$ (44)
in which, in order to regularize the integrals, the substitution ${\mathbf{p}}_{l}]\rightarrow[{\mathbf{p}}_{l}]+il\epsilon$ has been made. Using the Sochocki formula this can be then rewritten (assuming that $\epsilon>0$ - it will be seen that the result does not depend on the sign of $\epsilon$) in the form
${F^{(N)pp}\over V}=(-1)^{N-1}~\!{C_{0}^{N}\over N}\!\int_{\mathbf{q}}\sum_{n=1}^{N}\int_{{\mathbf{p}}_{n}}\!\!N_{--}^{{\mathbf{p}}_{n}}\left\{\prod_{j=1}^{n-1}\int_{{\mathbf{p}}_{j}}\!\left({\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]}-i\pi~\!\{{\mathbf{p}}_{j}\}~\!\delta([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])\right)\right$.
$\left.\times\prod_{j=n+1}^{N}\int_{{\mathbf{p}}_{j}}\!\left({\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]}+i\pi~\!\{{\mathbf{p}}_{j}\}~\!\delta([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])\right)\right\},$
in which the integrals of the factors $\{{\mathbf{p}}_{j}\}/([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])$ are understood in the Principal Value sense. The experience with the order $(k_{\rm F}a_{0})^{2}$ and $(k_{\rm F}a_{0})^{3}$ contributions teaches that removing divergences amounts simply to replacing $\{{\mathbf{p}}_{j}\}$ by $\{{\mathbf{p}}_{j}\}^{\rm sub}$ in the first terms of the integrands of the integrals over ${\mathbf{p}}_{j}$-s (but not in the delta-terms) and $C_{0}^{N}$ in front by $(C_{0}^{\rm R})^{N}$. This would be obvious had the dimensional regularization been used to handle ultraviolet divergences - by definition it sets the integrals like $\int_{\mathbf{p}}({\rm const})$ to zero and, as is well known (see e.g. [7]), $C_{0}=C_{0}^{\rm R}$ to all orders, if such a regularization is used. After the change of the variables ${\mathbf{p}_{n}}={\mathbf{k}}_{1}$, ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the order $(k_{\rm F}a_{0})^{N}$ particle-particle diagram contribution to $F$ can be then neatly written in the form
${F^{(N)pp}\over V}={C_{0}^{\rm R}\over N}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})\sum_{n=1}^{N}(L+iL_{\delta})^{n-1}(L-iL_{\delta})^{N-n}~\!.$ (45)
Summing the (finite) geometric series then gives
${F^{(N)pp}\over V}={C_{0}^{\rm R}\over N}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})~\!{(L+iL_{\delta})^{N}-(L-iL_{\delta})^{N}\over 2iL_{\delta}}~\!.$ (46)
[…]
${F^{pp}\over V}=C_{0}^{\rm R}\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})~\!{{\rm arctan}(L_{\delta}/(1-L))\over L_{\delta}}~\!.$ (47)
Refer to caption Refer to caption
Figure 3: The difference $(F(P)-F(0))^{pp}/N$ (in units $(3/5)\varepsilon_{\rm F}$) for $T=0$ and $T=0.2~\!T_{\rm F}\equiv 0.2~\!\varepsilon_{\rm F}/k_{\rm B}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values of the gas parameter $k_{\rm F}a_{0}$.
The zero temperature analog of this formula (i.e. representing the contribution of the particle-particle ring diagrams to the ground state energy density $E/V$) has been for $P=0$ first given by Kaiser [31] who in deriving it relied on combinatoric arguments. The formula which is the zero temperature analog of (47) for arbitrary polarization $P$ has been then written (by invoking the Kaiser’s reasoning) down in [18, 30] (see also [32]).
Refer to caption Refer to caption
Figure 4: As in Figure 3 but for $T=0.3~\!T_{\rm F}$ and $T=0.5~\!T_{\rm F}$.
The numerical procedure for evaluating the expression (30) described in details in [29] (the main trick is to construct - for given values of $t=k_{\rm B}T/\varepsilon_{\rm F}$ and $P$, which together, through (18), determine the chemical potentials - the interpolations of the functions of the parameter $\Delta$ into which the function (31) can be decomposed) can be used to evaluate also the expression (47). Figures 3 and 4 show for four different values of the temperature and several values of the gas parameter $k_{\rm F}a_{0}$ the difference $(F(P)-F(0))^{pp}/V$ (in units $(k^{3}_{\rm F}/3\pi^{2})(3/5)\varepsilon_{\rm F}$) obtained by adding to the zeroth order term (17) given explicitly by the formula (25) in [29] the contribution (47). In agreement with the result obtained in the last section of the work [18], one can observe that for $T=0$ the minimum of $(F(P)-F(0))^{pp}/V$ starts to move away from $P=0$ for $k_{\rm F}a_{0}=0.79$ indicating the continuous transition to the ordered state. As can be expected, with increasing temperature this critical value of the expansion parameter shifts towards larger values (0.85, 0.92 and 1.12 for $T=0.2~\!T_{\rm F}$, $0.3~\!T_{\rm F}$ and $0.5~\!T_{\rm F}$, respectively). One can also see that the minimum is back at $P=0$ for $k_{\rm F}a_{0}>0.96$ (at $T=0$) - this is the first order “reentrant” transition to the paramagnetic state observed in [20, 30] which is the consequence of the existence of the maximum of the energy density (at $T=0$) for $P=0$ treated as a function of $k_{\rm F}a_{0}$ shown by the dashed (blue) lines in the left panel of Figure 5. (From the physical point of view this reentrant transition is largely irrelevant as it occurs for the values of the gas parameter for which the formation of dimers prevails and the free energy of the metastable state is no longer physical.) This maximum of the energy density (indeed seen in experiments with cold gases [33]) which occurs close to the Feshbach resonance on its so-called BEC side (i.e. for large positive scattering length $a_{0}$) and the existence of which for higher temperatures (for which there is no phase transition) has been given a theoretical explaination (using a completely different approach) in [34] results here, as has been shown in [18, 20], from the appearance for $k_{\rm F}a_{0}>1.34$ of the simple pole in the “in-medium” particle-particle elastic scattering amplitude which can be interpreted as being due to the “in-medium” positive energy bound state of two fermions (of opposite spin projections). The dashed (blue) lines on the right panel of this figure and in Figures 6 and 7 illustrate how the contribution (47) to the free energy and its maximum change with the polarization and the temperature.
As the right panel of Figure 3 and Figure 4 show, with increasing temperature the reentrant transition occurs for higher ($1.06$ and $1.46$ for $T=0.2~\!T_{\rm F}$ and $T=0.3~\!T_{\rm F}$, respectively and yet higher for $T=0.5~\!T_{\rm F}$) values of the expansion parameter. The maximal depth of the minimum (at which $P\neq 0$) of $F$ first slightly decreases with the growing temperature (up to $T\approx 0.2~\!T_{\rm F}$) and then increases with it. Similarly $P=1$ is for temperatures up to $T\approx 0.2~\!T_{\rm F}$ reached only for $k_{\rm F}a_{0}$ values approaching the one at which the reentrant transition takes place but for higher temperatures it is reached well before it.
5. The particle-hole diagrams
At the order $(k_{\rm F}a_{0})^{3}$ to the free energy contributes also the second diagram shown in Figure 2. The corresponding analytical expression is given by the convolution
${F^{(3)ph}\over V}={C_{0}^{3}\over 3}~\!{1\over\beta}\sum_{l}\!\int_{\mathbf{q}}[B(\omega^{\rm B}_{l},{\mathbf{q}})]^{3}~\!,$ (48)
of three $B$-blocks which have the form [29]
$B(\omega_{l},{\mathbf{q}})=\int_{\mathbf{p}}{\{{\mathbf{p}}\}\over i\omega^{\rm B}_{l}-[{\mathbf{p}}]}~\!,$ (49)
analogous to (26) but with now different meaning of the symbols $\{{\mathbf{p}}\}$ and $\mathbf{p}$}]:
$N^{\mathbf{p}}_{+-}\equiv[1-n_{+}({\mathbf{q}}+{\mathbf{p}})]~\!n_{-}({\mathbf{p}})~\!,$
$N^{\mathbf{p}}_{-+}\equiv n_{+}({\mathbf{q}}+{\mathbf{p}})~\![1-n_{-}({\mathbf{p}})]~\!,$ (50)
$\{{\mathbf{p}}\}\equiv n_{+}({\mathbf{q}}+{\mathbf{p}})-n_{-}({\mathbf{p}})=-N^{\mathbf{p}}_{+-}(1-e^{\beta[{\mathbf{p}}]})~\!,$
$[{\mathbf{p}}]\equiv\varepsilon_{\mathbf{p}}-\tilde{\mu}_{-}^{(0)}-\varepsilon_{{\mathbf{q}}+{\mathbf{p}}}+\tilde{\mu}_{+}^{(0)}~\!,$
After performing in (48) the summation over the bosonic Matsubara frequencies and using (50) one arrives at
${F^{(3)ph}\over V}=-{C_{0}^{3}\over 3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\int_{{\mathbf{p}}_{3}}\!\left(N_{+-}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{\{{\mathbf{p}}_{3}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms}\right),$ (51)
where “two other terms” means terms in which the role of ${\mathbf{p}}_{1}]$ is played by ${\mathbf{p}}_{2}]$ and ${\mathbf{p}}_{3}]$. One can again make contact with the order $(k_{\rm F}a_{0})^{3}$ contribution of the particle-hole diagram of Figure 2 to the ground state energy density $E/V$ computed in [14, 15] where it was given as a sum of two functions $K_{1}$ and $K_{2}$ (Eq. (17) in [14]), by using the algebraic identity
$N_{+-}^{{\mathbf{p}}_{1}}~\!{N_{-+}^{{\mathbf{p}}_{2}}-N_{+-}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{N_{-+}^{{\mathbf{p}}_{3}}-N_{+-}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms}\phantom{aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa}$ (52)
$=N_{+-}^{{\mathbf{p}}_{1}}~\!{N_{-+}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{N_{-+}^{{\mathbf{p}}_{3}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+N_{-+}^{{\mathbf{p}}_{1}}~\!{N_{+-}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{N_{+-}^{{\mathbf{p}}_{3}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms},$
(it is in fact the identity (35) but written with different symbols). Using the symmetry of this expression allows to write the expression for $F^{(3)ph}/V$ as the sum of two terms which in the zero temperature limit reproduce the two terms, $K_{1}$ and $K_{2}$, which in [14, 15] represented the contribution of the second diagram of Figure 2 to $E/V$.
Refer to caption Refer to caption
Figure 5: Dependence on the gas parameter $k_{\rm F}a_{0}$ of the resumed contributions of the particle-particle diagrams given by the expressions (47) (dashed blue lines), of the particle-hole diagrams given by (60) (dotted red lines) and of their sum (solid green lines) for zero temperature and two values of the polarization $P$.
In order to properly treat the singularities in (51) we again make the substitutions ${\mathbf{p}}_{l}]\rightarrow[{\mathbf{p}}_{l}]+il\epsilon$ which allow to profit from the symmetry of the three terms of (51). One then, similarly as in the case of $F^{(3)pp}$, obtains
${F^{(3)ph}\over V}=-C_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!N_{+-}^{{\mathbf{p}}_{1}}\!\left[\left({\rm P}\!\int_{{\mathbf{p}}_{2}}{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\right)^{2}+{1\over 3}\left(i\pi\!\!\int_{{\mathbf{p}}_{2}}\!\{{\mathbf{p}}_{2}\}~\!\delta([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}])\right)^{2}\right].$~ (53)
Making now the changes of the variables: first ${\mathbf{p}}_{1}+{\mathbf{q}}={\mathbf{k}}_{1}$, ${\mathbf{p}}_{1}=-{\mathbf{k}}_{2}$, and then in the $n_{+}$ terms of $\{{\mathbf{p}}_{2}\}$ the change ${\mathbf{p}}_{2}+{\mathbf{k}}_{1}+{\mathbf{k}}_{2}=-{\mathbf{p}}$ and ${\mathbf{p}}_{2}={\mathbf{p}}$ in the $n_{-}$ terms of $\{{\mathbf{p}}_{2}\}$, and then taking explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the expression for $F^{(3)ph}/V$ can be written in the form
${F^{(3)ph}\over V}=-C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\![1-n_{+}({\mathbf{k}}_{1})]~\!n_{-}({\mathbf{k}}_{2})\left[K^{2}({\mathbf{k}}_{1},{\mathbf{k}}_{2})+{1\over 3}\left(iK_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})\right)^{2}\right],$ (54)
in which the dimensionless functions $K$ and $K_{\delta}$ are given by the integrals
[…]
$\left.+\int_{0}^{\infty}\!dp~\!p~\!n_{-}(p)\ln\!\left|{p-\Delta_{2}\over p+\Delta_{2}}\right|\right),$ (55)
$K_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})=\pi~\!{C_{0}^{\rm R}m_{f}\over(2\pi)^{2}\hbar^{2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}\left(\int_{|\Delta_{1}|}^{\infty}\!dp~\!p~\!n_{+}(p)-\int_{|\Delta_{2}|}^{\infty}\!dp~\!p~\!n_{-}(p)\right),$ (56)
in which
$\Delta_{1}\equiv{{\mathbf{k}}_{1}\!\cdot\!({\mathbf{k}}_{1}+{\mathbf{k}}_{2})\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!,\phantom{aaa}\Delta_{2}\equiv{{\mathbf{k}}_{2}\!\cdot\!({\mathbf{k}}_{1}+{\mathbf{k}}_{2})\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!.$ (57)
Again, the limits of the integrals in $K_{\delta}$ are determined by the condition that the zeroes of the arguments of the Dirac deltas treated as functions of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lie between 1-1 and +1+1. And again the function $K_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can be written down in a closed analytical form - see below.
Refer to caption Refer to caption
Figure 6: As in Figure 5 but for $T=0.2~\!T_{\rm F}$.
The expression (54) is finite although this is not immediately obvious: while the integrals defining the functions $K$ and $K_{\delta}$ are clearly (ultraviolet) finite owing to the presence of the distribution functions $n_{+}(p)$ and $n_{-}(p)$, the integral over $d^{3}{\mathbf{k}}_{1}$ has no such an exponentially suppressing factor. In addition, the presence of the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of the functions $K$ and $K_{\delta}$ seem to imply the potential presence of a singularity in the integral over the directions of ${\mathbf{k}}_{1}$ (for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$). To prove the finiteness of (54) we will first analyze the behavior of the difference $K^{2}-K_{\delta}^{2}/3$ in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$. To this end it is helpful to split the functions $K$ and $K_{\delta}$ into $K_{+}+K_{-}$ and $K_{\delta+}+K_{\delta-}$ (the subscripts $\pm$ corresponds to the $n_{\pm}$ distribution functions). The hardest to see is the convergence of the integral over $d^{3}{\mathbf{k}}_{1}$ involving the factor $K_{-}^{2}-K^{2}_{\delta-}/3$. The dangerous term $n_{-}({\mathbf{k}}_{2})[K_{-}^{2}-K^{2}_{\delta-}/3]$ comes from the term
$\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\int_{{\mathbf{p}}_{3}}\left\{{n_{-}({\mathbf{p}}_{1})~\!n_{-}({\mathbf{p}}_{2})~\!n_{-}({\mathbf{p}}_{3})\over([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}])([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}])}+{\rm two~other~terms}\right\},$
in (51). This, however is algebraically zero (just as the integrand of (36)) and, therefore, the factor $K_{-}^{2}-K^{2}_{\delta-}/3$ must be zero too. As any inaccuracy in the numerical evaluation of the integrals (55) and (56) could lead to a nonzero difference $K_{-}^{2}-K^{2}_{\delta-}/3$ and, therefore, to a (fake) nonconvergence of the integration over $d^{3}{\mathbf{k}}_{1}$, in the term with unity arising from $1-n_{+}({\mathbf{k}}_{1})]$ we simply replace $K^{2}-K^{2}_{\delta}/3$ by $K_{+}^{2}+2K_{+}K_{-}-(K_{\delta+}^{2}+2K_{\delta+}K_{\delta-})/3$. The rest of the terms are separately integrable in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$:
$K_{\delta+}\propto{1\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!\ln\!\left(1+e^{-\beta(\hbar^{2}\Delta_{1}^{2}/2m-\tilde{\mu}^{(0)}_{+})}\right)\approx{1\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!e^{-\beta(\hbar^{2}\Delta_{1}^{2}/2m-\tilde{\mu}^{(0)}_{+})}~\!,$
because $\Delta^{2}_{1}$ grows like ${\mathbf{k}}_{1}^{2}$ as $|{\mathbf{k}}_{1}|\rightarrow\infty$. This secures the convergence of the integrals of the terms $K_{\delta+}^{2}$ and $2K_{\delta+}K_{\delta-}$. Similarly, it can be estimated that the integral over $p$ in $K_{+}$ behaves as $1/\Delta_{1}$ when $|{\mathbf{k}}_{1}|\rightarrow\infty$. Since each of the $K_{\pm}$ functions has the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of it, the term $(K_{+})^{2}$ behaves for $|{\mathbf{k}}_{1}|\rightarrow\infty$ as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\cdot{\mathbf{k}}_{2})^{2}$ and this secures the convergence of the integration over $d^{3}{\mathbf{k}}_{1}$. The term $2K_{+}K_{-}$, instead, behaves only as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\!\cdot\!{\mathbf{k}}_{2})|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, but the integration over the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$ kills the term of order $1/|{\mathbf{k}}_{1}|^{3}$ and the remaining terms are already integrable. This only power-like suppression of the integration of the term $K_{+}^{2}+2K_{+}K_{-}$ makes, however, numerical evaluation of the particle-hole contribution to the free energy more difficult and, therefore, potentially less accurate than the evaluation of the particle-particle one. As to the potentially singular factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, one should first notice that the original expressions (48) and (49) as well as similar formulae giving the contributions of the $N$-th order particle-hole rings do not contain such singularities. They are due to the symmetrizations needed to arrive at the final formulae and must, therefore, cancel out (like the spurious singularities of the integrand in (36)) even if it is not directly evident. In the third order finitness of (54) can be seen as follows: since the integrals in the definitions of the $K_{\pm}$ and $K_{\delta\pm}$ functions are finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$, the singularities of (54) have essentially the form $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{2}^{2}+2|{\mathbf{k}}_{1}||{\mathbf{k}}_{2}|\xi)$ where $\xi$ is the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$. In the third order contribution considered here, after integration over $\xi$ they give rise to terms $\ln\!||{\mathbf{k}}_{1}|-|{\mathbf{k}}_{2}||$ which are integrable. [^5] In fact for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$ the factors $\Delta_{1}$ and $\Delta_{2}$ behave as $\sqrt{1+\xi}$ and it can be checked (numerically) that the integrals in the functions $K_{+}$ and $K_{-}$ vanish then as $\sqrt{1+\xi}$, so these functions have therefore finite limits. The singularities reside only in the terms involving the functions $K_{\delta\pm}$. In the numerical evaluation of the integrals in (54) and in (60) it is however sufficient to impose a cutoff $|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|>\kappa$ and check that the results stabilize as $\kappa$ approaches zero. Thus the expression (54) is finite and we have checked that evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) it reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-hole diagram of Figure 2 obtained in [14, 15].
Refer to caption Refer to caption
Figure 7: As in Figure 5 but for $T=0.5~\!T_{\rm F}$.
Using the same tricks as previously the contribution to the free energy of the infinite series of Feynman diagrams composed of $N$-fold rings of the particle-hole $B$-blocks of Figure 1 can be summed in a closed form. The order $(k_{\rm F}a_{0})^{N}$ term of this series is
${F^{(N)ph}\over V}=-(-1)^{N+1}~\!{C_{0}^{N}\over N}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\dots\!\int_{{\mathbf{p}}_{N}}\sum_{n=1}^{N}N_{+-}^{{\mathbf{p}}_{n}}\left(\prod_{j\neq n}^{N}{\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]+i(n-j)\epsilon}\right),$ (58)
(apart from the extra minus originating from the minus sign in the identity $\{{\mathbf{p}}\}=-N^{\mathbf{p}}_{+-}(1-e^{\beta[{\mathbf{p}}]})$, the origin of the rest of the prefactor is the same as in the case of $F^{(N)pp}$) and, repeating the steps one arrives at the formal sum
$\sum_{N=1}^{\infty}{F^{(N)ph}\over V}=-~\!C_{0}^{\rm R}\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\![1-n_{+}({\mathbf{k}}_{1})]~\!n_{-}({\mathbf{k}}_{2})~\!{{\rm arctan}(K_{\delta}/(1-K))\over K_{\delta}}~\!.$ (59)
From this formal sum one has to subtract two first terms of the series: there is no order $C_{0}^{\rm R}$ particle-hole diagram and the order $(C_{0}^{\rm R})^{2}$ term is already (recall that the order $C_{0}^{2}$ contribution of the diagram shown in Figure 1 can be written either as a convolution of two $A$-blocks or of two $B$-blocks shown in the same Figure), in the properly renormalized form, included in (47). Thus the final form of the ressumed contributions of the particle-hole ring diagrams is
${F^{ph}\over V}=-C_{0}^{\rm R}\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\![1-n_{+}({\mathbf{k}}_{1})]~\!n_{-}({\mathbf{k}}_{2})\left[{{\rm arctan}(K_{\delta}/(1-K))\over K_{\delta}}-1-K\right].$ (60)
One has to comment again on the finiteness of the expression (60). The singularities related to the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ are now harmless because the function $K$ is, as remarked, finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$ and the singular function $K_{\delta}$ is now in the denominator and under the arctan function. As to the ultraviolet finiteness of (60), the integral of the factor $K$ explicitly subtracted in the square brackets is ultraviolet divergent being simply equivalent to the divergent expression (28). To ensure proper cancelation of this term in (60) for large $|{\mathbf{k}}_{1}|$ and/or $|{\mathbf{k}}_{2}|$ we expand the arctan function in $K_{\delta}$ and $K$ up to the sixth order (we have checked that taking more terms of the expansion does not change the result appreciably). The ultraviolet divergent terms linear in $K$ then disappear and, moreover, the expansion allows to implement the discussed trick with replacing $K^{2}-K_{\delta}^{2}/3$ by $K^{2}_{+}+2K_{+}K_{-}-(K^{2}_{\delta+}+2K_{\delta+}K_{\delta-})$.
In [15] we have compared the order $(k_{\rm F}a_{0})^{3}$ contributions to the ground state energy (i.e. to the free energy $F$ for zero temperature) of the particle-particle and of the particle-hole diagrams (shown in Figure 2) in the entire range of the polarization $P$ and found that the second one is not much smaller than the first one. In Figures 5, 6 and 7 we compare the magnitudes of the resummed contributions (47) and (60) to the free energy as functions of $k_{\rm F}a_{0}$ for three different temperatures and two representative values of the system’s polarization $P$. We also plot the sum of these two contributions. It again follows that for $k_{\rm F}a_{0}\sim 1$ the resummed contribution of the particle-hole diagrams is not much smaller than that of the particle-particle ones and is of the opposite sign. It is also clear that in this most important region the sum of the two contributions is significantly distorted compared to the resummed contribution of the particle-particle diagrams. This raises the question what impact the resummed contribution of the particle-hole ring diagrams has on the results reported in the papers [18, 20, 30].
Refer to caption Refer to caption
Figure 8: The difference $(F(P)-F(0))^{pp+ph}/N$ (in units $(3/5)\varepsilon_{\rm F}$ for $T=0$ and $T=0.2~\!T_{\rm F}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values of the gas parameter $k_{\rm F}a_{0}$.
Figures 8 and 9, analogous to Figures 3 and 4 illustrate the consequences of adding the resummed contribution (60) of the particle-hole ring diagrams to the free energy. The result is dramatic: the phase transition to the ordered state discussed in [18, 20, 30] completely disappears! Thus, even if the selection of the particle-particle ring diagrams as giving the dominant contribution can be (partially) justified by invoking the arguments, given in [35], based on using $1/2^{D/2}$ where $D$ is the number of space dimensions as the expansion parameter, they in practice do not turn out to be really dominant: the contribution of other subsets of diagrams (the number of such subsets beginning at a given order of the expansion grows with the order number) can, as our results show, change qualitatively the behavior of the thermodynamical potentials of the system of fermions close to the Feshbach resonance.
7. Conclusions
We have applied the systematic thermal (imaginary time) perturbative expansion to the effective (low energy) field theory to compute the free energy of the gas of interacting (nonrelativistic) spin $1/2$ fermions for arbitrary values of the gas polarization and temperatures not exceeding the Fermi temperature. We have shown how to circumvent the technical problem which previously prevented us from immediately extending such a computation beyond the second order in the gas parameter $k_{\rm F}a_{0}$ and have given explicit formulae for the order $(k_{\rm F}a_{0})^{3}$ contributions to the system’s free energy. It turned out that the analytical part of this computations is more transparent and easier than the corresponding direct computations of the ground state energy based on the formula (14) which gives only the zero temperature limit of the results obtained with the help of the thermal expansion. (Of course, numerical evaluation of the resulting expressions for a nonzero temperature is considerably more involved than for $T=0$).
Refer to caption Refer to caption
Figure 9: As in Figure 8 but for $T=0.3~\!T_{\rm F}$ and $T=0.5~\!T_{\rm F}$.
To obtain the complete order $(k_{\rm F}R)^{3}$ contribution to the free energy of the gas of fermions interacting through a truly repulsive spin-independent two-body potential (characterized by a length scale $R$) one would have to add the contributions arising from the operators of lower length dimension in the effective theory interaction term because such potentials naturally give the $p$-wave scattering length $a_{1}$ and the $s$-wave effective range $r_{0}$ comparable to the $s$-wave scattering length $a_{0}$. Instead of doing this, in this work we have profited from the simple structure of the contributions of the particle-particle and particle-hole ring diagrams and managed to give simple formulae for their contributions to the free energy resummed to all orders in the gas parameter $k_{\rm F}a_{0}$. These formulae apply therefore rather to cold gases of fermionic atoms (interacting through attractive poentials) close to the Feshbach resonance where their $s$-wave scattering length $a_{0}$ is made positive and much larger than the remaining scattering lengths and effective radii. Using these formulae we have first checked that including only the contributions of the particle-particle rings we reproduce for zero temperatures all the results obtained in [18]. In particular we confirm that in this approximation at zero temperature the transition to the ordered phase occurs for $k_{\rm F}a_{0}=0.79$ and that it is continuous. These results seem to agree well with the results of the dedicated Monte Carlo computations. Our formula would, however, allow to study also the thermal profile of the transition.
However we have found that the phase transition to the ordered state completely disappears after including into the free energy the resummed contribution of the particle-hole ring diagrams - the minimum of the free energy is always for zero polarization. This may indicate at least that the agreement of the critical value of the gas parameter $k_{\rm F}a_{0}$ found in the papers [18, 20, 30] with its value obtained from the Monte Carlo simulations (done with attractive potentials tuned so that $a_{0}$ is positive and large) is just accidental. It may however also indicate that there is indeed no transition (in the metastable state) to the spin ordered phase if the true interaction is attractive (but $a_{0}$ is made positive and large by approaching a Feshbach resonance). The vanishing at zero temperature of the inverse spin susceptibility ($1/\chi\propto(\partial^{2}F/\partial P^{2})_{P=0}$) for some value of the gas parameter found in the Monte Carlo simulations (which in the case of attractive potentials are not as clean as in the repulsive case because of the need to exclude - in the finite volume, that is without taking the thermodynamic limit - an overlap with the true ground state of the considered system of atoms) [21] from which one infers its existence and its continuous character would then be misleading (the second order approximation to $F$ predicts analytically vanishing of the inverse spin susceptibility at $k_{\rm F}a_{0}=1.058$ whereas the transition is first order and occurs for $k_{\rm F}a_{0}=1.054$ [18]). If this hypothesis is true it would provide yet another (in addition to the formation of atomic dimers) reason for the failure to simulate the intinerant ferromagnetism with the help of cold atoms. In an case it is clear that more theoretical studies are needed to clarify the situation.