### v2

# Free energy of the gas of spin-1/2 fermions beyond the second order and the Stoner phase transition

Abstract
Applying the previously developed systematic thermal (imaginary time) perturbative expansion to the relevant effective field theory we compute the free energy $F$ of the diluted gas of (nonrelativistic) spin $1/2$ fermions interacting through a spin-independent repulsive two-body potential as a function of the numbers $N_{+}$ and $N_{-}$ of spin up and spin down fermions (i.e. as a function of the system’s polarization) and the temperature $T$. We give the complete order $(k_{\rm F}a_{0})^{3}$ ($k_{\rm F}$ is the Fermi wave vector and $a_{0}$ is the $s$-wave scattering length characterizing the interaction potential) contribution to $F$. We also extend the computation beyond a fixed order by resumming to all orders in the parameter $k_{\rm F}a_{0}$ the contributions to $F$ of two infinite sets of Feynman diagrams: the so-called particle-particle rings and the particle-hole rings. We find that including the second one of these two contributions has a dramatic consequence for the transition of the system from the paramagnetic to the ferromagnetic phase (the so called Stoner phase transition): in this approximation the phase transition simply disappears. This result does not contradict the expectation that a transition to the magnetically ordered state should occur in truly repulsive systems. The $p$-wave and higher scattering lengths, as well as other parameters chacterizing the interaction potential, are in such systems generally of the same order of magnitude as $a_{0}$ and contributions depending on them should be, therefore, also included in $F$. Our results may, however, have implications for the search of the itinerant ferromagnetism of cold atomic gases in which large $a_{0}$, much larger than all other parameters, is artificially created by exploiting the physics of the Feshbach resonance.
Keywords: Diluted gas of interacting fermions, effective field theory, itinerant ferromagnetism, phase transitions.
1. Introduction
In recent years some progress has been achieved in the computation of equilibrium zero temperature properties of the gas of $N$ (nonrelativistic) fermions interacting with each other through a spin independent repulsive potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$. It was mainly related to the application to this classic [1, 2, 3, 4, 5, 6] many-body quantum mechanics and statistical physics problem of general methods (see e.g. [7]) of the effective field theory. In this approach, initiated in the seminal paper [8] (see also [9]), the original spatially nonlocal potential $V_{\rm pot}$ assumed to be characterized by a length scale $R$ is replaced by the a priori infinite series of local interactions (written here using the standard second quantization formalism - see e.g. Ref. [6] - with $\hat{\psi}_{\pm}$ being the field operators of spin up and down fermions)
$\hat{V}_{\rm int}=C_{0}\!\int\!d^{3}{\mathbf{x}}~\!(\hat{\psi}_{+}^{\dagger}\hat{\psi}_{+})(\hat{\psi}_{-}^{\dagger}\hat{\psi}_{-})+\hat{V}_{\rm int}^{(C_{2})}+\hat{V}_{\rm int}^{(C_{2}^{\prime})}+\dots,$ (1)
of decreasing length dimension. The coefficients (couplings) $C_{0}$, $C_{2}$, $C_{2}^{\prime},\dots$ of the interaction (1) can be then directly determined in terms of the quantities - the scattering lengths $a_{0},a_{1},\dots$ and the effective radii $r_{0},\dots$ - parametrizing the general expansion (in powers of the relative momentum) of the amplitude of the elastic scattering of two particles. Trading the (bare) couplings of (1) for these measurable quantities which characterize the underlying potential $V_{\rm pot}$ of the binary interactions also has the effect of removing UV infinities engendered by the locality of (1). The simplifications brought in by this approach allowed first to easily reproduce [8] those terms of the perturbative expansion of the (ground-state) energy density $E/V$ of the system of spin $s$ fermions with equal densities of different spin projections which in the past were obtained by more traditional (and requiring considerably more work) methods of many-body quantum mechanics (such as the ones based on Green’s functions or on the Galitskii approach) [6] and to extend [10] this computation to the fourth order in the systematic, organized by the power counting rules [8], expansion in powers (in higher orders modified also by logarithms) of the dimensionless product $k_{\rm F}R$ of the system’s overall Fermi wave vector
$k_{\rm F}=(6\pi^{2}n/g_{s})^{1/3}~\!,$ (2)
(here $n=N/V$ is the overall density of the gas of spin $s$ fermions and $g_{s}=2s+1$) and the characteristic length $R$. The same approach allowed also to compute up to order $(k_{\rm F}a_{0})^{2}$ the ground state energy of the system of spin $1/2$ fermions for different densities of fermions with the up and down spin projections [11] recovering the old result of Kanno [12] (obtained analytically for the hard core interaction potential) and to easily extend it [13] to fermions having spin $s$ greater than 1/2 (and, thus, more possible spin projections). Finally, in Refs. [14, 15] and slightly later in [16] this computation has been extended up to the third order in the systematic expansion in powers of $k_{\rm F}R$.
These technical achievements allowed to investigate more quantitatively the phase transition, also called the Stoner transition, to the ordered state in which the densities $N_{+}/V$ and $N_{-}/V$ of fermions of opposite spin projections in the case of $s=1/2$ are not equal and the system exhibits a nonzero polarization $P\equiv(N_{+}-N_{-})/N$. According to the standard qualitative argument based on the positivity of the repulsive interaction energy and the Pauli exclusion principle, for sufficiently strong repulsion and/or sufficiently high overall density, such an ordered state should be energetically (at zero temperature) favored over the state with equal densities. In this regime the system should, therefore, exhibit the property called itinerant ferromagnetism. It has been found that the clear first order character of this transition predicted by the second order of the perturbative expansion (in agreement with the general quantitative arguments given long ago in Ref. [17]) - at sufficiently large values of the “gas parameter” $k_{\rm F}a_{0}$ two symmetric minima of $E(P)$ are formed away from $P=0$ separated from the one at $P=0$ by a finite barrier and at $(k_{\rm F}a_{0})_{\rm cr}=1.054$ they become the global minima ($|P_{\rm cr}|=0.58$) - gets significantly weakened ($|P_{\rm cr}|$ shifts significantly toward $P=0$ and the barrier becomes much lower) if the complete order $(k_{\rm F}a_{0})^{3}$ contribution is taken into account [14, 15]. This seemed to be in line with the results of the work [18] in which a certain class of contributions to $E/V$ (arising from the so-called particle-particle ring diagrams) has been resummed to all orders in $k_{\rm F}a_{0}$ finding that in this approximation the transition is continuous - the new minima start to continuously move away from $P=0$ as the gas parameter crosses some critical value $(k_{\rm F}a_{0})_{\rm cr}\approx 0.8$ ($0.858$ if only 1 hole-hole N1N-1 particle-particle parts of the $N$-th order particle-particle ring diagrams and $0.79$ if the complete $N$-th order particle-particle ring diagrams are resummed; see also Ref. [19] for a refinement of this approach). It should be added that the predictions of Ref. [18], which qualitatively seem to be supported by the results of a different approach [20], agree at first sight quite well with the results obtained with the help of the Monte Carlo simulations [21].
However if all scattering lengths $a_{\ell}$ and effective radii $r_{\ell}$ are of the same order of magnitude $\sim\!R$, the complete (according to the power counting rules [8] which apply, strictly speaking, only to such a case) order $(k_{\rm F}R)^{3}$ contribution to the energy density depends also on $a_{1}$ and $r_{0}$ and it has been shown in Ref. [16] that in this case the character of the phase transition (at zero temperature) predicted by this approximation depends on the relative magnitudes and signs of the parameters $a_{1}$, $r_{1}$ and $a_{0}>0$ (that is, on the more detailed characteristics of the underlying potential $V_{\rm pot}$). The computation (such as the one done in Ref. [18]) in which to all orders resummed are (some) contributions depending only on powers of $k_{\rm F}a_{0}$ correspond rather to the situation encountered in physics of dilute atomic gases in which the $s$-wave scattering length $a_{0}$ (it can be of either sign) is made positive and very large compared to the remaining parameters ($|a_{0}|\gg R\sim|a_{1}|,|r_{1}|,\dots$) by exploiting properties of the Feshbach resonances (see e.g. Ref. [22]). In this case however the underlying interaction of fermions (atoms) is attractive and the scattering length $a_{0}$ is positive in the regime in which bound states composed of two fermions of opposite spins can form. The true ground state of the system is then very different from the one of nointeracting atoms (they may not be adiabatically connected to one another in the thermodynamic limit implicit in the field theory approach) which is used in perturbative computations performed within the effective field theory approach. In view of that the energy density $E/V$ obtained perturbatively by expanding around the ground-state of the system of noninteracting fermions most probably is the energy density of the metastable (from the thermodynamic point of view) state in which a real gas of cold fermionic atoms finds itself just after switching the external magnetic field so that the $s$-wave scattering length $a_{0}$ becomes positive and large. However, although the dependence on the polarization of this energy density indicates that the transition to the ferromegnetic state (which in principle could occur in a metastable state [23]) should occur, a too short lifetime of the gas metastable state may prevent its observation in real systems. Therefore the failure of the attempts to observe such a transition [24, 25, 26] in experiments is usually attributed to the too rapid formation (at very low temperatures at which these experiments were carried out) of atomic dimers (bosons). It should however be remarked that in the situation in which the underlying interaction is attractive (despite giving rise to a positive $s$-wave scattering length) the mentioned qualitative argument for the occurrence of the transition no longer applies and the expectation that the transition should occur is mainly based on the textbook mean field correction to the energy density [4, 5] (equivalent to the first order correction of the perturbative expansion) which depends only linearly on the $s$-wave scattering length $a_{0}$.
Computation of the ground-state energy density $E/V$ as a function of the densities of fermions with different spin projections allows only to investigate the equilibrium properties of the system of interacting fermions at zero (or a very low) temperature. It is however of interest to determine its behavior also at nonzero temperatures. (In the context of the physics of atomic gases it is physically clear that at higher temperatures the mentioned formation of atomic dimers should be less important.) This requires computing one of the thermodynamic potentials of the system of interacting fermions. So far such a computation of the free energy $F$ has been done [27] only up to the second order, i.e. up to terms of order $(k_{\rm F}a_{0})^{2}$, using the old-fashioned thermal perturbation theory (see e.g. Ref. [28], paragraph 32) based on the ordinary second order Rayleigh - Schrödinger perturbative expression for energy levels entering the statistical sum. In Ref. [29] we have applied to the effective field theory (1) the standard thermal perturbative expansion (exploiting the imaginary time formalism - see e.g. Ref. [6]) which is routinely used to compute thermodynamic characteristics of much more complex systems like e.g. the quark-gluon plasma. Using it we recovered this second order expression for $F$ and reproduced the thermal characteristics of the Stoner phase transition it predicts (however, pointing out problems - not discussed in [27] - with accurate numerical determination of the critical values of the polarization), but have encountered a technical problem which prevented us from immediately extending the computation to higher orders. Here we show how this problem can be resolved and applying the developed systematic thermal expansion to the first term of the effective field theory interaction (1) we derive the formulas allowing to compute numerically the complete order $(k_{\rm F}a_{0})^{3}$ corrections to the free energy $F(T,V,N_{+},N_{-})$. With additional work, including the contributions of the next two terms of (1) it would be possible to compute the complete order $(k_{\rm F}R)^{3}$ correction to the free energy.
In this paper, instead of completing the order $(k_{\rm F}R)^{3}$ correction to the free energy (which would allow to investigate in more detail the thermal profile of the Stoner phase transition to the ordered state induced by truly repulsive spin-independent two-body potetials $V_{\rm pot}$ which necessarily give rise to the parameters $a_{1}$ and $r_{0}$ of comparable magnitude to that of $a_{0}$), we profit from the possibility provided by the simpler structure of the terms of the expansion generated by the imaginary time formalism and resum to all orders in $k_{\rm F}a_{0}$ not only the contributions to the temperature-dependent free energy $F$ of the particle-particle ring diagrams (done for zero temperature in Refs. [18, 19, 30]) but also of the particle-hole ring diagrams. The thermodynamic potential obtained in this way is most probably the free energy of the system of fermionic atoms in a metastable state on the BEC side of the Feshbach resonance and can therefore serve to study its properties.
Including in the free energy $F(T,V,N,P)$ only the resummed contribution of the particle-particle ring diagrams we recover for $T=0$ the results of the works [18, 30] and can study how they are modified at nonzero temperatures not exceeding the Fermi temperature (we expect that the results obtained within the effective field theory should be valid for temperatures in this range). However we find that the inclusion of the contribution of the resummed particle-hole diagrams changes the situation drastically: the phase transition to the ordered state simply disappears (the minimum of $F$ is at $P=0$ for all values of the parameter $k_{\rm F}a_{0}$ and all temperatures). This is a somewhat surprising result and we comment on its possible meaning for the problem of itinerant ferromagnetism in system of cold gases in the Conclusions.
2. The formalism
The natural statistical formalism in which to compute equilibrium properties of the gas of fermions the interaction of which preserves their spins and therefore the numbers $N_{\pm}$ of spin up and spin down particles, is the grand canonical ensemble with two independent chemical potentials $\mu_{\pm}$. The relevant statistical operator is then
$\hat{\rho}={1\over\Xi_{\rm stat}}~\!e^{-\beta\hat{K}}~\!$, (3)
where $\beta\equiv 1/k_{\rm B}T$, with $T$ the temperature and $k_{\rm B}$ the Boltzmann constant, and
$\hat{K}=\hat{H}_{0}-\mu_{+}\hat{N}_{+}-\mu_{-}\hat{N}_{-}+\hat{V}_{\rm int}\equiv\hat{K}_{0}+\hat{V}_{\rm int}~\!.$ (4)
[…]
$\hat{K}_{0}=\sum_{\sigma=\pm}\!\int_{\mathbf{p}}(\varepsilon_{\mathbf{p}}-\mu_{\sigma})a^{\dagger}_{\sigma}({\mathbf{p}})a_{\sigma}({\mathbf{p}})~\!,$ (5)
with $\varepsilon_{\mathbf{p}}=\hbar^{2}{\mathbf{p}}^{2}/2m_{f}$. The standard systematic thermodynamical perturbative expansion [6, 29] gives the potential $\Omega$ in the form of the series in powers of the interaction $\hat{V}_{\rm int}$
$\Omega=\Omega^{(0)}-{1\over\beta}\sum_{N=1}^{\infty}{(-1)^{N}\over N!}\!\int_{0}^{\beta}\!d\tau_{N}\dots\!\int_{0}^{\beta}\!d\tau_{1}~\!{\rm Tr}\!\left(\hat{\rho}^{(0)}{\rm T}_{\tau}[\hat{V}_{\rm int}^{I}(\tau_{N})\dots\hat{V}_{\rm int}^{I}(\tau_{1})]\right)^{\rm con}.$ (6)
Here $\hat{V}_{\rm int}^{I}(\tau)=e^{\tau\hat{K}_{0}}\hat{V}_{\rm int}e^{-\tau\hat{K}_{0}}$ is the interaction operator in the (imaginary time) interaction picture, ${\rm T}_{\tau}$ is the chronological ordering and $\hat{\rho}^{(0)}$ is the statistical operator of the noninteracting system. The superscript “con” means that only connected contributions (Feynman diagrams) should be taken into account. The first term in (6) is the textbook [4, 5] grand thermodynamical potential of the system of nointeracting fermions
$\Omega^{(0)}(T,V,\mu_{+},\mu_{-})=-{V\over\beta}\sum_{\sigma=\pm}\int_{\mathbf{p}}\!\ln\!\left(1+e^{-\beta(\varepsilon_{\mathbf{p}}-\mu_{\sigma})}\right).$ (7)
Owing to the thermal analog of the Wick formula (see e.g. Ref. [6]) computation of the successive terms $\Omega^{(N)}$ with $N\geq 1$ of the expansion (6) reduces to drawing all possible connected Feynman diagrams with $N$ interaction vertices arising from $\hat{V}_{\rm int}$ joined by oriented lines and integrating over positions $\mathbf{x}$} and “times” $\tau$ labeling these vertices the corresponding products of the free thermal propagators
$-{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}(\tau_{2}-\tau_{1},{\mathbf{x}}_{2}-{\mathbf{x}}_{1})={1\over\beta}\sum_{n\in{\mathbb{Z}}}\int_{\mathbf{p}}e^{-i\omega_{n}^{\rm F}(\tau_{2}-\tau_{1})}~\!e^{i{\mathbf{p}}\cdot({\mathbf{x}}_{2}-{\mathbf{x}}_{1})}\left(-\tilde{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}(\omega_{n}^{\rm F},{\mathbf{p}})\right),$ (8)
the Fourier transform $\tilde{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}$ of which have the form [6]
$-\tilde{\cal G}^{(0)}_{\sigma_{2}\sigma_{1}}(\omega_{n}^{\rm F},{\mathbf{p}})={-\delta_{\sigma_{2}\sigma_{1}}\over i\omega_{n}^{\rm F}-(\varepsilon_{\mathbf{p}}-\mu_{\sigma})}~\!,$ (9)
associated with (oriented) lines connecting vertices of each diagram. The resulting “momentum” space Feynman rules are almost identical with the ordinary ones except that integrations over frequencies (energies) are replaced by summations over the (fermionic) Matsubara frequencies $\omega_{n}^{\rm F}=(\pi/\beta)(2n+1)$, $n\in{\mathbb{Z}}$.
Applying this formalism with the interaction operator $\hat{V}_{\rm int}$ given by the first term of (1) one finds that the order $C_{0}$ term of the expansion (6) is simply given by
$\Omega^{(1)}=C_{0}V{\cal G}_{++}(0,{\mathbf{0}})~\!{\cal G}_{--}(0,{\mathbf{0}})~\!,$ (10)
with
${\cal G}_{\pm\pm}(0,{\mathbf{0}})=\int_{\mathbf{p}}\left[1+e^{\beta(\varepsilon_{\mathbf{p}}-\mu_{\pm})}\right]^{-1}~\!.$ (11)
Higher-order contributions to the potential $\Omega$ can also be systematically computed. If $\hat{V}_{\rm int}$ in (4) were the true, spatially nonlocal, two-body interaction [corresponding to a two-body potential $V_{\rm pot}(|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}|)$, where ${\mathbf{x}}_{i}$ are the positions of fermions], the successive terms of the expansion (6) would be (ultraviolet) finite. If $\hat{V}_{\rm int}$ is the local interaction (1) of the effective theory, the successive terms of the expansion (6) involve ultraviolet divergences and have to be regularized. As in our previous works we will employ for this purpose the cutoff $\Lambda$ on the wave vectors of virtual particles. Finite (in the limit $\Lambda\rightarrow\infty$) contributions to the potential $\Omega$ (a physical quantity) are then obtained by systematically expressing the (bare) couplings of $\hat{V}_{\rm int}$ in terms of other measurable (physical) quantities. As it is customary, and in line (at least when the gas is very diluted) with the physical intuition that properties of the gas are mainly determined by elastic two-body collisions of its constituents, one expresses the couplings $C_{0}$, $C_{2}$, etc. of (1) in terms of the measurable quantities related to such a scattering process, namely in terms of the scattering lengths $a_{0}$, $a_{1}$, effective ranges $r_{0}$, etc. [8]. In this work we will only need to express the coupling $C_{0}$ in this way; the relevant formula obtained by matching the amplitude of the elastic fermion-fermion scattering computed perturbatively using the first term of the interaction (1) onto the general form of the same amplitude parameterized by $a_{0}$, $a_{1},\dots$ and $r_{0},\dots,$ reads [11, 14, 10]
$C_{0}={4\pi\hbar^{2}\over m_{f}}~\!a_{0}\left(1+{2\over\pi}~\!a_{0}\Lambda+{4\over\pi^{2}}~\!a^{2}_{0}\Lambda^{2}\dots\right)\equiv C_{0}^{\rm R}\left(1+{2\over\pi}~\!a_{0}\Lambda+{4\over\pi^{2}}~\!a^{2}_{0}\Lambda^{2}\dots\right).$ (12)
From the thermodynamic point of view much more convenient to work with than the potential $\Omega$ is the free energy $F$ which canonically depends on the variables $T$, $V$ and the particle numbers $N_{\pm}$ which are given by the derivatives
$N_{\pm}=-(\partial\Omega/\partial\mu_{\pm})_{T,V}~\!.$ (13)
In principle, in each order of the expansion (6) to construct the free energy one should invert the relations (13) to obtain the chemical potentials as functions of the particle numbers $N_{+}$ and $N_{-}$ and insert these in the formula $F=\Omega+\mu_{+}N_{+}+\mu_{-}N_{-}$. Thus the values of the chemical potentials $\mu_{+}$ and $\mu_{-}$ change with each successive order of the expansion and the procedure of constructing the free energy looks at first sight rather cumbersome. It turns out, however, that in the systematic expansion this procedure simplifies considerably: it amounts in effect to using the chemical potentials $\mu_{\pm}^{(0)}$ determined by inverting the formula (13) with $\Omega$ replaced by $\Omega^{(0)}$ given by (7) and omitting in the expansion (6) those diagrams which give vanishing contribution in computing the corrections $\Delta E$ to the ground state energy $E=E^{(0)}+\Delta E$ of the system of interacting particles using the ordinary Dyson expansion of the formula [8, 10, 11, 14, 15] ($T$ stands here for time, not for the temperature, and ${\rm T}_{t}$ means the ordinary chronological ordering)
$\Delta E=\lim_{T\rightarrow\infty}{i\hbar\over T}~\!\langle 0|{\rm T}_{t}\exp\!\left({1\over i\hbar}\!\int_{-T/2}^{T/2}\!dt~\!V_{\rm int}^{I}(t)\right)\!|0\rangle~\!,$ (14)
in which $|0\rangle$ is the ground state of the noninteracting system of $N=N_{+}+N_{-}$ fermions. In the case of the interaction proportional to $C_{0}$ this has been explicitly demonstrated in [29] up to the third order of the perturbative expansion. This prescription is obviously consistent with the fact that in the zero temperature limit the corrections to the free energy obtained from the thermodynamic expansion should go over into the corresponding corrections to the ground state energy given by (14).
[…]
can also be easily taken into account in this formalism by including it in the free Hamiltonian $\hat{H}_{0}$ which amounts to shifting the chemical potentials $\mu_{\pm}\rightarrow\tilde{\mu}_{\pm}\equiv\mu_{\pm}\pm{\cal H}$ in $\hat{K}_{0}$ given by (4). The free energy is then given as the series
$F(T,V,{\cal H},N_{+},N_{-})=F^{(0)}+F^{(1)}+F^{(2)}+\dots,$ (16)
in which
$F^{(0)}(T,V,{\cal H},N_{+},N_{-})=\Omega^{(0)}(T,N,\tilde{\mu}_{+}^{(0)},\tilde{\mu}_{-}^{(0)})+(\tilde{\mu}_{+}^{(0)}-{\cal H})N_{+}+(\tilde{\mu}_{+}^{(0)}+{\cal H})N_{-}~\!,$ (17)
and $F^{(N)}(T,V,{\cal H},N_{+},N_{-})=\Omega^{(N)}(T,N,\tilde{\mu}_{+}^{(0)},\tilde{\mu}_{-}^{(0)})$ for $N=1,2,\dots,$ where, as explained above, in computing $\Omega^{(1)}$, $\Omega^{(2)}$ etc. one should take into account only those diagrams of the expansion (6) which give nonzero contributions to $\Delta E$. The (shifted) chemical potentials $\tilde{\mu}_{+}^{(0)}$, $\tilde{\mu}_{-}^{(0)}$ are given by
$\tilde{\nu}_{\pm}^{(0)}\equiv\tilde{\mu}_{\pm}^{(0)}/k_{\rm B}T=f^{-1}\!\left((1\pm P)\left({\varepsilon_{\rm F}(n)\over k_{\rm B}T}\right)^{3/2}\right),$ (18)
where $\varepsilon_{\rm F}(n)\equiv\hbar^{2}k_{\rm F}^{2}/2m_{f}$ and $f^{-1}(x)$ is the inverse of the monotonic function (mapping $\mathbb{R}$} onto $\mathbb{R}$$_{+}$) defined by the integral
$f(\nu)={3\over 2}\!\int_{0}^{\infty}\!d\xi~\!{\xi^{1/2}\over 1+e^{\xi-\nu}}~\!.$ (19)
If the computation of $F$ is restricted to the order $C^{\rm R}_{0}$ (i.e. to the order $k_{\rm F}a_{0}$) correction given, as follows from the formulated prescription and the result (10) and (11), by
$F^{(1)}=VC_{0}(N_{+}/V)(N_{-}/V)~\!,$ (20)
the condition for the minimum of $F$ with respect to $N_{+}$ and $N_{-}$ (at fixed $N_{+}+N_{-}=N$) which determines the system’s polarization $P$ takes the form ($t\equiv k_{\rm B}T/\varepsilon_{\rm F}$, $h\equiv{\cal H}/\varepsilon_{\rm F}$)
${8\over 3\pi}~\!(k_{\rm F}a_{0})~\!P+2h=t\left[f^{-1}\!\left({1+P\over t^{3/2}}\right)-f^{-1}\!\left({1-P\over t^{3/2}}\right)\right].$ (21)
If the asymptotic expansion
$f^{-1}(x)=x^{2/3}\left[1-(\pi^{2}/12)x^{-4/3}-(\pi^{4}/80)x^{-8/3}-(247\pi^{6}/25920)x^{-4}+\dots\right],$ (22)
valid for $x\gg 1$ [obtained by inverting the Sommerfeld expansion [28] of the function $f(\nu)$] is used, the formula (21) reproduces the textbook [4, 5] low temperature equilibrium condition (equivalent to the condition $\mu_{+}=\mu_{-}$) and leads to the well known prediction that the Stoner phase transition to the ordered state is continuous with divergent magnetic susceptibility characterized by the critical exponent $\gamma=1$ and a finite discontinuity of the heat capacity. (In fact, this continuous character of the transition is accidental: in the same approximation the transition is of first order in the system of spin $s>1/2$ fermions and/or if the space dimension is different than 3.) If the correction $F^{(2)}$ is included, the transition becomes first order, at least at sufficiently low temperatures [27, 11, 13] in agreement with the arguments given in the past in [17]. However the computation of the complete order $(k_{\rm F}a_{0})^{3}$ correction to the ground state energy $E$ performed in [14, 15, 16] showed that the first order character of the transition at zero temperature, very clear in the second order approximation, is significantly weakened, at least as long as the contributions proportional to $k^{3}_{F}a^{2}_{0}r_{0}$ and $(k_{F}a_{1})^{3}$ (which, if the underlying interaction potential $V_{\rm pot}$ is “natural”, i.e. if all $a_{\ell}$, $r_{\ell}$, etc. are of the same order of magnitude, are of the same, $(k_{\rm F}R)^{3}$, order as the $(k_{\rm F}a_{0})^{3}$ correction) are not taken into account [16]. Below we extend the existing computations in two ways: we compute the complete (proportional to $(C_{0}^{\rm R})^{3}$, i.e. to $(k_{\rm F}a_{0})^{3}$) temperature dependent third order corrections to the free energy $F$ and, moreover, we show how to perform the resummation of two infinite subsets of temperature dependent corrections to $F$ of which the first one is the finite temperature generalization of the subset of diagrams taken into account in (the last section of) ref. [18].
3. The order $(k_{\rm F}a_{0})^{2}$ and the order $(k_{\rm F}a_{0})^{3}$ particle-particle corrections to $F$
We begin by recalling the computation of the order $(C_{0}^{\rm R})^{2}$ term $F^{(2)}$ performed in Ref. [29]. In agreement with the formulated prescription it is given by the single Feynman diagram shown in figure 1. The corresponding analytical expression can be obtained by convoluting either two $A$-blocks or two $B$-blocks shown in the same figure:
$F^{(2)}=-{1\over 2}~\!C_{0}^{2}V~\!{1\over\beta}\sum_{l\in\mathbb{Z}}\!\int_{\mathbf{q}}[A(\omega_{l}^{B},\mathbf{q})]^{2}=-{1\over 2}~\!C_{0}^{2}V~\!{1\over\beta}\sum_{l\in\mathbb{Z}}\!\int_{\mathbf{q}}[B(\omega_{l}^{B},\mathbf{q})]^{2}.$ (23)
To make the formulas resulting from convoluting $A$-blocks more transparent it will be convenient to introduce the following notation:
$N_{--}^{\mathbf{p}}\equiv n_{+}({\mathbf{p}})~\!n_{-}({\mathbf{q}}-{\mathbf{p}})~\!,$
$N_{++}^{\mathbf{p}}\equiv[1-n_{+}({\mathbf{p}})]~\![1-n_{-}({\mathbf{q}}-{\mathbf{p}})]~\!,$
$n_{\pm}({\mathbf{p}})=\left[1+\exp\{\beta(\varepsilon_{\mathbf{p}}-\tilde{\mu}_{\pm}^{(0)})\}\right]^{-1},$ (24)
$\{{\mathbf{p}}\}\equiv n_{+}({\mathbf{p}})+n_{-}({\mathbf{q}}-{\mathbf{p}})-1~\!,$
$[{\mathbf{p}}]\equiv\varepsilon_{\mathbf{p}}-\tilde{\mu}_{+}^{(0)}+\varepsilon_{{\mathbf{q}}-{\mathbf{p}}}-\tilde{\mu}_{-}^{(0)}~\!.$
At zero temperature $N_{--}^{\mathbf{p}}$ and $N_{++}^{\mathbf{p}}$ reduce respectively to $\theta(p_{{\rm F}+}-|{\mathbf{p}}|)\theta(p_{{\rm F}-}-|{\mathbf{q}}-{\mathbf{p}}|)$ and $\theta(|{\mathbf{p}}|-p_{{\rm F}+})\theta(|{\mathbf{q}}-{\mathbf{p}}|-p_{{\rm F}-})$, hence the subscripts -- and ++++. It is also easy to check that
$\{{\mathbf{p}}\}=N_{--}^{\mathbf{p}}-N_{++}^{\mathbf{p}}=N_{--}^{\mathbf{p}}\left(1-e^{\beta[{\mathbf{p}}]}\right).$ (25)
The form of the distribution functions $n_{\pm}({\mathbf{p}})$ plays the role only in the second of these two identities.
Figure 1: The order $C_{0}^{2}$ diagram contributing to the thermodynamic potential $F$ (or to the ground-state energy $E$) of the gas of spin $1/2$ fermions and two “elementary” one-loop diagrams (AA- and $B$-blocks) out of which the second order and those higher order (in the $C_{0}$ coupling) contributions that are taken into account in this work are composed. Solid and dashed lines denote propagators of fermions with the spin projections ++ and -, respectively.
In the introduced notation the $A$-block (obtained in Ref. [29]) takes the form
$A(\omega_{l}^{B},\mathbf{q})=\int_{\mathbf{p}}\!{\{{\mathbf{p}}\}\over i\omega_{l}^{B}-[{\mathbf{p}}]}~\!,$ (26)
After the sum in (23) over the bosonic Matsubara frequencies $\omega^{B}_{l}=(\pi/\beta)l$ is performed using the standard formulas [6, 29] one gets
${F^{(2)}\over V}=-{1\over 2}~\!C_{0}^{2}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!{\{{\mathbf{p}}_{1}\}\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\left({1\over 1-e^{\beta[{\mathbf{p}}_{1}]}}-{1\over 1-e^{\beta[{\mathbf{p}}_{2}]}}\right).$ (27)
Since the two terms are formally identical (after making in the integrals in one of the terms the interchange ${\mathbf{p}}_{1}\leftrightarrow{\mathbf{p}}_{2}$), one arrives, using (25), at the final form of $F^{(2)}/V$:
${F^{(2)}\over V}=C_{0}^{2}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!N_{--}^{{\mathbf{p}}_{1}}~\!{1\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}-C_{0}^{2}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!,$ (28)
in which $\{{\mathbf{p}}\}^{\rm sub}\equiv\{{\mathbf{p}}\}+1=n_{+}({\mathbf{p}})+n_{-}({\mathbf{q}}-{\mathbf{p}})$ and the integrals should be understood in the principal value sense. Notice that the denominators ${\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]$ do not depend on the chemical potentials. This profiting from the symmetry of the two terms of (27), seemingly not problematic, has indeed no consequences here but, as will be shown, in higher orders if applied blindly would lead to incorrect results.
The first term in (28) is divergent. The change of the variables ${\mathbf{q}}=2{\mathbf{s}}$, ${\mathbf{p}}_{1}={\mathbf{s}}-{\mathbf{t}}_{1}$, and ${\mathbf{p}}_{2}={\mathbf{s}}-{\mathbf{t}}_{2}$ (the Jacobian equals 8) makes the innermost integral elementary and allows to write it in the form
${16\pi^{2}\hbar^{4}\over m^{2}_{f}}~\!a^{2}_{0}\left(1+{4\over\pi}~\!a_{0}\Lambda+\dots\right)\!\int_{\mathbf{s}}\!\int_{{\mathbf{t}}_{1}}\!8~\!n_{+}({\mathbf{s}}-{\mathbf{t}}_{1})~\!n_{-}({\mathbf{s}}+{\mathbf{t}}_{1})~\!{m_{f}\over 2\pi^{2}\hbar^{2}}\left(-\Lambda+{{\mathbf{t}}^{2}_{1}\over\Lambda}+\dots\right),$
after using (12).
Expressing $C_{0}$ similarly in the second term of the formula (28) and in $F^{(1)}$ given by (20), one finds that the divergent terms of order $a_{0}^{2}\Lambda$ cancel out and
${F^{(1)}+F^{(2)}\over V}=C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})-(C_{0}^{\rm R})^{2}\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}$
$-{16\hbar^{2}\over\pi m_{f}}~\!a_{0}^{3}\!\int_{\mathbf{s}}\!\int_{{\mathbf{t}}_{1}}8~\!n_{+}({\mathbf{s}}-{\mathbf{t}}_{1})~\!n_{-}({\mathbf{s}}+{\mathbf{t}}_{1})~\!(\Lambda^{2}-2{\mathbf{t}}_{1}^{2})\phantom{aaaaaaaaa}$~\! (29)
$-{64\pi\hbar^{4}\over m_{f}^{2}}~\!a_{0}^{3}\Lambda\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\ N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}+{\cal O}(1/\Lambda)~\!.\phantom{aaaaaaaa}$~\!
The first two terms constitute the complete, finite contribution to $F/V$ up to the order $(C_{0}^{\rm R})^{2}$; the remaining terms are formally of higher order and can be considered only after including other third and higher order contributions.
In Ref. [29] it has been found that it is convenient to evaluate (the finite part of) $F^{(2)}/V$ by substituting ${\mathbf{p}}_{1}={\mathbf{k}}_{1}$, and ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$, ${\mathbf{p}}_{2}={\mathbf{p}}$ (the Jacobian is 1), replacing (by another change of the integration variable) $n_{-}({\mathbf{k}}_{1}+{\mathbf{k}}_{2}-{\mathbf{p}})$ with $n_{-}({\mathbf{p}})$ and then performing explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$. This allows to represent the order $(k_{\rm F}a_{0})^{2}$ contribution to $F$ in the form
${F^{(2)}\over V}=C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})~\!L({\mathbf{k}}_{1},{\mathbf{k}}_{2})~\!.$ (30)
The (dimensionless) function $L({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by the single integral
$L({\mathbf{k}}_{1},{\mathbf{k}}_{2})=-{C_{0}^{\rm R}m_{f}\over(2\pi)^{2}\hbar^{2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}\!\int_{0}^{\infty}\!dp~\!p~\![n_{+}(p)+n_{-}(p)]\ln\!\left|{(p-\Delta_{+})(p-\Delta_{-})\over(p+\Delta_{+})(p+\Delta_{-})}\right|,$ (31)
in which
$\Delta_{\pm}={1\over 2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|\pm{1\over 2}|{\mathbf{k}}_{1}-{\mathbf{k}}_{2}|~\!.$ (32)
In [29] we have checked that in the zero temperature limit, in which the Fermi distribution functions $n_{+}({\mathbf{p}})$ and $n_{-}({\mathbf{p}})$ are replaced by the step functions $\theta(p_{{\rm F}+}-|{\mathbf{p}}|)$ and $\theta(p_{{\rm F}-}-|{\mathbf{p}}|)$, this formula reproduces numerically the second order correction to the ground state energy computed first (analytically) by Kanno [12] and then recovered (semi-analytically) in several works (e.g. in [11, 13]) for all values of the polarization $P$. We have also analyzed the free energy $F$ with the corrections $F^{(1)}$ and $F^{(2)}$ included and recovered, up to uncertainties following from the finite precision of the (rather complicated) numerical evaluation of the relevant multiple integrals the characteristics of the phase transition to the ordered state (for temperatures $T<\varepsilon_{\rm F}/k_{\rm B}$) first obtained in [27].
Figure 2: The particle-particle and particle-hole diagrams contributing in the order $C^{3}_{0}$ to the thermodynamic potential $F$ (or to the ground-state energy $E$).
If only the interaction term proportional to $C_{0}$ in (1) is taken into account, there are two Feynman diagrams contributing to the free energy $F$ in the third order. The first one, shown in the left panel of Fig. 2, is termed the particle-particle ring diagram. Its contribution $F^{(3)pp}$ is given by the convolution of three $A$-blocks
${F^{(3)pp}\over V}={1\over 3}~\!C_{0}^{3}~\!{1\over\beta}\sum_{l}\int_{\mathbf{q}}\![A(\omega_{l}^{\rm B},\mathbf{q})]^{3}~\!.$ (33)
After decomposing the product of three $A$-blocks into simple fractions, performing the summation over the bosonic Matsubara frequencies $\omega_{l}^{\rm B}$ and then using the identities (25) one arrives at
${F^{(3)pp}\over V}={1\over 3}~\!C_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\int_{{\mathbf{p}}_{3}}\!\left(N_{--}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{\{{\mathbf{p}}_{3}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms}\right),$ (34)
where “two other terms” means the terms in which the role of ${\mathbf{p}}_{1}$ is played by ${\mathbf{p}}_{2}$ and ${\mathbf{p}}_{3}$. It is good at this point to make a contact with the contribution of this third order particle-particle ring diagram to the ground state energy density $E/V$ obtained in Refs. [14, 15] (and in Refs. [16]) to which the expression (34) should reduce in the zero temperature limit, that is when the Fermi distribution functions are replaced by the theta functions. The contribution to $E/V$ of the particle-particle diagram was given in Refs. [14, 15] by two terms (Eq. (21) in Ref. [15]) whereas here it is given by the single term (three seemingly identical terms). The equivalence of the two approaches is ensured by the algebraic, i.e. independent of the precise forms of $N_{--}^{\mathbf{p}}$ and $\mathbf{p}$}] (recall that $\{{\mathbf{p}}\}=N_{--}^{\mathbf{p}}-N_{++}^{\mathbf{p}}$), identity which results from the symmetrization
$N_{--}^{{\mathbf{p}}_{1}}~\!{N_{--}^{\mathbf{p}_{2}}-N_{++}^{\mathbf{p}_{2}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{2}]}~\!{N_{--}^{\mathbf{p}_{3}}-N_{++}^{\mathbf{p}_{3}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{3}]}+{\rm two~other~terms}\phantom{aaaaaaaaaaaaaaaaaaaaaaaaaaaaa}$ (35)
$=\left(N_{--}^{\mathbf{p}_{1}}~\!{N_{++}^{\mathbf{p}_{2}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{2}]}~\!{N_{++}^{\mathbf{p}_{3}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{3}]}+N_{++}^{\mathbf{p}_{1}}~\!{N_{--}^{\mathbf{p}_{2}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{2}]}~\!{N_{--}^{\mathbf{p}_{3}}\over[\mathbf{p}_{1}]-[\mathbf{p}_{3}]}\right)+{\rm two~other~terms}.$
After using the symmetry, i.e. taking only the content of the bracket and multiplying it by 3, it allows to rewrite the expression for $F^{(3)pp}/V$ as the sum of two terms which in the $T=0$ limit precisely reduce to the two terms, $G_{1}$ and $G_{2}$, which in Refs. [14, 15] contributed to $E/V$.
Naively, as all three terms of (34) seem also identical, one is tempted to compute only one of them and multiply the result by three. $F^{(3)pp}/V$ would be in this way given by a single fourfold integral. This, as we have found, would lead to an incorrect result which in the zero temperature limit would not agree with the contribution of the particle-particle diagram to $E/V$ (this is precisely the problem that did not allow us to immediately extend the computation reported in Ref. [29]).
To understand the problem it is instructive to consider the triple integral
$\int_{0}^{1}\!dx\!\int_{0}^{1}\!dy\!\int_{0}^{1}\!dz\left({1\over(x-y)(x-z)}+{1\over(y-x)(y-z)}+{1\over(z-x)(z-y)}\right).$ (36)
The integrand is algebraically zero and the result of the integration should be zero too. Yet the integrand has (spurious) singularities and the integrals in (36), similarly as the ones encountered in the computation of $F^{(3)}$, should be understood in the principal value sense. If one naively says that the integrals of the three terms are equal and evaluates only one of them (multiplying it by 3) one will get
$3\int_{0}^{1}\!dx\!\left({\rm P}\!\int_{0}^{1}\!dy~\!{1\over x-y}\right)^{2}=3\int_{0}^{1}\!dx~\!\ln^{2}{1-x\over x}=3~\!{\pi^{2}\over 3}.$
The correct result (zero) is obtained if one first regularizes the integrand of (36) by setting $x\rightarrow x+i\epsilon$, $y\rightarrow y+2i\epsilon$, $z\rightarrow z+3i\epsilon$ (the sign of $\epsilon$ is irrelevant; the integrand is still algebraically zero but its singularities are now off the integration axes). It is then straightforward to find that the application of the Sochocki formula $1/(x\pm i0)=P(1/x)\mp i\pi\delta(x)$ to the regularized integral (36) leads to (the terms linear in the Dirac deltas neatly cancel out)
$3\int_{0}^{1}\!dx\!\left({\rm P}\!\int_{0}^{1}\!dy~\!{1\over x-y}\right)^{2}+\int_{0}^{1}\!dx\!\left(i\pi\!\int_{0}^{1}\!dy~\!\delta(x-y)\right)^{2}=0~\!.$
If the same procedure is applied to (34) one gets
${F^{(3)pp}\over V}=C_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!~N_{--}^{{\mathbf{p}}_{1}}\!\left[\left({\rm P}\!\int_{{\mathbf{p}}_{2}}~\!{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\right)^{2}+{1\over 3}\left(i\pi\!\int_{{\mathbf{p}}_{2}}\!\{{\mathbf{p}}_{2}\}~\!\delta([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}])\right)^{2}\right].$ (37)
One can now check that in the sum $F^{(1)}+F^{(2)}+F^{(3)pp}$ all the divergences [up to the order $(k_{\rm F}a_{0})^{3}$] cancel out (as will be seen, the contribution $F^{(3)ph}$ of the other diagram of Figure 2, which completes the order $(k_{\rm F}a_{0})^{3}$ contribution to $F^{(3)}$, is finite; this also follows from the computations of the order $(k_{\rm F}a_{0})^{3}$ corrections to $E/V$ performed in Refs. [14, 15]). Writing $\{{\mathbf{p}}_{i}\}=-1+\{{\mathbf{p}}_{i}\}^{\rm sub}$ in the first term in the square bracket in (37) allows to single out the divergent part of $F^{(3)}/V$. It is given by (to this order one can set in (37) $C_{0}=C_{0}^{\rm R}$; we also suppress the symbol $P$ of the principal value)
${F^{(3)pp}_{\rm div}\over V}={64\pi^{3}\hbar^{6}\over m_{f}^{3}}~\!a_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}N_{--}^{{\mathbf{p}}_{1}}\!\left(\int_{{\mathbf{p}}_{2}}~\!{1\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\right)^{2}\phantom{aaaaaaaaaaa}$
$-2~\!{64\pi^{3}\hbar^{6}\over m_{f}^{3}}~\!a_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}N_{--}^{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}~\!{\{{\mathbf{p}}_{2}\}^{\rm sub}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\int_{{\mathbf{p}}_{3}}~\!{1\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}~\!.$ (38)
Making now in the first term the change of the variables ${\mathbf{q}}=2{\mathbf{s}}$, ${\mathbf{p}}_{1}={\mathbf{s}}-{\mathbf{t}}_{1}$, ${\mathbf{p}}_{2}={\mathbf{s}}-{\mathbf{t}}_{2}$ (the Jacobian is 8) and performing the innermost integral (over $d^{3}{\mathbf{t}}_{2}$) one finds that it precisely cancels the entire middle line of (29). Moreover, after making similar changes of the variables in the last line of (29) and in the last term of $F^{(3)pp}_{\rm div}/V$ they too mutually cancel out. The remaining contribution of the left diagram of Figure 2 is, therefore, given by (37) with $\{{\mathbf{p}}_{2}\}$ in the first term (but not in the second one!) replaced by $\{{\mathbf{p}}_{2}\}^{\rm sub}$ and $C_{0}$ replaced by $C_{0}^{\rm R}$. Making as previously the change ${\mathbf{p}}_{1}={\mathbf{k}}_{1}$, ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$, ${\mathbf{p}}_{2}={\mathbf{p}}$ of the integration variables one can represent the contribution of the particle-particle order $(k_{\rm F}a_{0})^{3}$ diagram to $F/V$ in the form
${F^{(3)pp}_{\rm fin}\over V}=C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{1}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})\left[L^{2}({\mathbf{k}}_{1},{\mathbf{k}}_{2})+{1\over 3}\left(iL_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})\right)^{2}\right],$ (39)
where the function $L({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by (31) while the dimensionless function $L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by the finite integral
$L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})=\pi~\!{C_{0}^{\rm R}m_{f}\over(2\pi)^{2}\hbar^{2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}\!\int_{p_{\rm min}}^{p_{\rm max}}\!dp~\!p~\![n_{+}(p)+n_{-}(p)-1]~\!,$ (40)
in which $p_{\rm min}=|\Delta_{-}|$ and $p_{\rm max}=\Delta_{+}$ are determined by the condition that the zero of the argument of the Dirac delta treated as a function of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lies between 1-1 and +1+1. Since $\varepsilon_{\mathbf{p}}$ depends on $p^{2}$, the function $L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can, of course, be obtained in a closed analytic form. The expression (39) is finite (as its superscript indicates) the ultraviolet convergence [^3] The singularity introduced by the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|^{2}$ is superficial because for ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}={\mathbf{0}}$ vanishes also the logarithm in (31) while in (40) $p_{\rm min}=p_{\rm max}$. of the integrations being secured by the exponential suppression provided by the Fermi distribution functions $n_{+}({\mathbf{k}}_{1})$ and $n_{-}({\mathbf{k}}_{2})$. We have also checked that the expression (39) evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-particle diagram of Figure 2 obtained in Refs. [14, 15].
4. Resummation of the contributions of the particle-particle diagrams
It turns out that the contribution to the free energy $F$ of the infinite series of Feynman diagrams composed of $N$-fold rings of the particle-particle $A$-blocks of Figure 1 can be summed in a closed form. Consider first the order $(k_{\rm F}a_{0})^{N}$ term of this series [the factor $(-1)^{N+1}$ is the same as in (6) - there are as many rearrangements of the $\hat{\psi}_{+}$ operators as of the $\hat{\psi}_{-}$ ones; the factor $1/N!$ in (6) is reduced to $1/N$ as there are (N1)!(N-1)! identical diagrams]:
${F^{(N)pp}\over V}=(-1)^{N+1}~\!{C_{0}^{N}\over N}~\!{1\over\beta}\sum_{l}\int_{\mathbf{q}}[A(\omega_{l}^{\rm B},{\mathbf{q}})]^{N}~\!.$ (41)
Decomposing the product of the integrands of the NN AA-blocks (26) using the identity
$\prod_{i=1}^{N}{1\over x-a_{i}}=\sum_{n=1}^{N}\left(\prod_{j\neq n}^{N}{1\over a_{n}-a_{j}}\right){1\over x-a_{n}}~\!,$ (42)
and performing then the summation over the Matsubara frequencies one gets the integrand of the $(N+1)$-fold integral in the form
$\{{\mathbf{p}}_{1}\}\dots\{{\mathbf{p}}_{N}\}\sum_{n=1}^{N}\left(\prod_{j\neq n}^{N}{1\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]}\right){1\over 1-e^{\beta[{\mathbf{p}}_{n}]}}~\!,$ (43)
and finally, after using the relations (25), $F^{(N)pp}/V$ takes the form
${F^{(N)pp}\over V}=(-1)^{N+1}~\!{C_{0}^{N}\over N}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\dots\!\int_{{\mathbf{p}}_{N}}\sum_{n=1}^{N}N_{--}^{{\mathbf{p}}_{n}}\left(\prod_{j\neq n}^{N}{\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]+i(n-j)\epsilon}\right),$ (44)
in which, in order to regularize the integrals, the substitution ${\mathbf{p}}_{l}]\rightarrow[{\mathbf{p}}_{l}]+il\epsilon$ has been made. Using the Sochocki formula this can be then rewritten (assuming that $\epsilon>0$ – it will be seen that the result does not depend on the sign of $\epsilon$) in the form
${F^{(N)pp}\over V}=(-1)^{N-1}~\!{C_{0}^{N}\over N}\!\int_{\mathbf{q}}\sum_{n=1}^{N}\int_{{\mathbf{p}}_{n}}\!\!N_{--}^{{\mathbf{p}}_{n}}\left\{\prod_{j=1}^{n-1}\int_{{\mathbf{p}}_{j}}\!\left({\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]}-i\pi~\!\{{\mathbf{p}}_{j}\}~\!\delta([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])\right)\right$.
$\left.\times\prod_{j=n+1}^{N}\int_{{\mathbf{p}}_{j}}\!\left({\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]}+i\pi~\!\{{\mathbf{p}}_{j}\}~\!\delta([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])\right)\right\},$
in which the integrals of the factors $\{{\mathbf{p}}_{j}\}/([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])$ are understood in the principal value sense. The experience with the order $(k_{\rm F}a_{0})^{2}$ and $(k_{\rm F}a_{0})^{3}$ contributions teaches that removing divergences amounts simply to replacing $\{{\mathbf{p}}_{j}\}$ by $\{{\mathbf{p}}_{j}\}^{\rm sub}$ in the first terms of the integrands of the integrals over ${\mathbf{p}}_{j}$-s (but not in the delta-terms) and $C_{0}^{N}$ in front by $(C_{0}^{\rm R})^{N}$. This would be obvious had the dimensional regularization been used to handle ultraviolet divergences –
Refer to caption Refer to caption
Figure 3: The difference $(F(P)-F(0))^{pp}/N$ (in units $(3/5)\varepsilon_{\rm F}$) for $T=0$ and $T=0.2~\!T_{\rm F}\equiv 0.2~\!\varepsilon_{\rm F}/k_{\rm B}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values (indicated in the panes) of the gas parameter $k_{\rm F}a_{0}$.
by definition it sets the integrals like $\int_{\mathbf{p}}({\rm const})$ to zero and, as is well known (see e.g. Ref. [8]), $C_{0}=C_{0}^{\rm R}$ to all orders, if such a regularization is used. After the change of the variables ${\mathbf{p}_{n}}={\mathbf{k}}_{1}$, and ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the order $(k_{\rm F}a_{0})^{N}$ particle-particle diagram contribution to $F$ can be neatly written in the form
${F^{(N)pp}\over V}={C_{0}^{\rm R}\over N}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})\sum_{n=1}^{N}(L+iL_{\delta})^{n-1}(L-iL_{\delta})^{N-n}~\!.$ (45)
Summing the (finite) geometric series then gives
${F^{(N)pp}\over V}={C_{0}^{\rm R}\over N}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})~\!{(L+iL_{\delta})^{N}-(L-iL_{\delta})^{N}\over 2iL_{\delta}}~\!.$ (46)
[…]
${F^{pp}\over V}=C_{0}^{\rm R}\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\!n_{+}({\mathbf{k}}_{1})~\!n_{-}({\mathbf{k}}_{2})~\!{{\rm arctan}(L_{\delta}/(1-L))\over L_{\delta}}~\!.$ (47)
The zero temperature analog of this formula (i.e. representing the contribution of the particle-particle ring diagrams to the gas energy density $E/V$) has been for $P=0$ first given by Kaiser [31] who in deriving it relied on combinatoric arguments. The formula that is the zero temperature analog of (47) for arbitrary polarization $P$ has then been written (by invoking the Kaiser’s reasoning) down in Refs. [18, 30] (see also Ref. [32]).
Refer to caption Refer to caption
Figure 4: As in Figure 3 but for $T=0.3~\!T_{\rm F}$ and $T=0.5~\!T_{\rm F}$.
The numerical procedure for evaluating the expression (30) described in detail in Ref. [29] [the main trick is to construct – for given values of $t=k_{\rm B}T/\varepsilon_{\rm F}$ and $P$, which together, through (18), determine the chemical potentials – the interpolations of the functions of the parameter $\Delta$ into which the function (31) can be decomposed] can be used to evaluate also the expression (47). Figures 3 and 4 show for four different values of the temperature and several values of the gas parameter $k_{\rm F}a_{0}$ the difference $(F(P)-F(0))^{pp}/V$ [in units $(k^{3}_{\rm F}/3\pi^{2})(3/5)\varepsilon_{\rm F}$] obtained by adding to the zeroth order term (17) given explicitly by the formula (25) in Ref. [29] the contribution (47). In agreement with the result obtained in the last section of the work [18], one can observe that for $T=0$ the minimum of $(F(P)-F(0))^{pp}/V$ starts to move away from $P=0$ for $k_{\rm F}a_{0}=0.79$ indicating the continuous transition to the ordered state. As can be expected, with increasing temperature this critical value of the expansion parameter shifts toward larger values (0.85, 0.92 and 1.12 for $T=0.2~\!T_{\rm F}$, $0.3~\!T_{\rm F}$ and $0.5~\!T_{\rm F}$, respectively) of the gas parameter $k_{\rm F}a_{0}$. One can also see that the minimum is back at $P=0$ for $k_{\rm F}a_{0}>0.96$ (at $T=0$) – this is the first order “reentrant” transition to the paramagnetic state observed in Refs. [19, 30] which is the consequence of the existence of the maximum of the energy density (at $T=0$) for $P=0$ treated as a function of $k_{\rm F}a_{0}$ shown by the dashed (blue) lines in the left panel of Figure 5. (From the physical point of view this reentrant transition is largely irrelevant as it occurs for the values of the gas parameter for which the formation of dimers prevails and the free energy of the metastable state is no longer physical.) This maximum of the energy density (indeed seen in experiments with cold gases [33]), which occurs close to the Feshbach resonance on its so-called BEC side (i.e. for large positive scattering length $a_{0}$) and the existence of which for higher temperatures (for which there is no phase transition) has been given a theoretical explanation (using a completely different approach) in Ref. [34], results here, as has been shown in Refs. [18, 19], from the appearance for $k_{\rm F}a_{0}>1.34$ of the simple pole in the “in-medium” particle-particle elastic scattering amplitude which can be interpreted as being due to the existence of the “in-medium” positive energy bound state of two fermions (of opposite spin projections). The dashed (blue) lines on the right panel of this figure and in Figs. 6 and 7 illustrate how the contribution (47) to the free energy and its maximum change with the polarization and the temperature.
As the right panel of Fig. 3 and Fig. 4 show, with increasing temperature the reentrant transition occurs for higher ($1.06$ and $1.46$ for $T=0.2~\!T_{\rm F}$ and $T=0.3~\!T_{\rm F}$, respectively and yet higher for $T=0.5~\!T_{\rm F}$) values of the expansion parameter. The maximal depth of the minimum (at which $P\neq 0$) of $F$ first slightly decreases with the rising temperature (up to $T\approx 0.2~\!T_{\rm F}$) and then increases with it. Similarly $P=1$ is for temperatures up to $T\approx 0.2~\!T_{\rm F}$ reached only for $k_{\rm F}a_{0}$ values approaching the one at which the reentrant transition takes place but for higher temperatures it is reached well before it.
5. The particle-hole diagrams
At the order $(k_{\rm F}a_{0})^{3}$ to the free energy contributes also the second diagram shown in Figure 2. The corresponding analytical expression is given by the convolution
${F^{(3)ph}\over V}={C_{0}^{3}\over 3}~\!{1\over\beta}\sum_{l}\!\int_{\mathbf{q}}[B(\omega^{\rm B}_{l},{\mathbf{q}})]^{3}~\!,$ (48)
of three $B$-blocks which have the form [29]
$B(\omega_{l},{\mathbf{q}})=\int_{\mathbf{p}}{\{{\mathbf{p}}\}\over i\omega^{\rm B}_{l}-[{\mathbf{p}}]}~\!,$ (49)
analogous to (26) but now with a different meaning of the symbols $\{{\mathbf{p}}\}$ and $\mathbf{p}$}]:
$N^{\mathbf{p}}_{+-}\equiv[1-n_{+}({\mathbf{q}}+{\mathbf{p}})]~\!n_{-}({\mathbf{p}})~\!,$
$N^{\mathbf{p}}_{-+}\equiv n_{+}({\mathbf{q}}+{\mathbf{p}})~\![1-n_{-}({\mathbf{p}})]~\!,$ (50)
$\{{\mathbf{p}}\}\equiv n_{+}({\mathbf{q}}+{\mathbf{p}})-n_{-}({\mathbf{p}})=-N^{\mathbf{p}}_{+-}(1-e^{\beta[{\mathbf{p}}]})~\!,$
$[{\mathbf{p}}]\equiv\varepsilon_{\mathbf{p}}-\tilde{\mu}_{-}^{(0)}-\varepsilon_{{\mathbf{q}}+{\mathbf{p}}}+\tilde{\mu}_{+}^{(0)}~\!.$
After performing in (48) the summation over the bosonic Matsubara frequencies and using (50) one arrives at
${F^{(3)ph}\over V}=-{C_{0}^{3}\over 3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\int_{{\mathbf{p}}_{3}}\!\left(N_{+-}^{{\mathbf{p}}_{1}}~\!{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{\{{\mathbf{p}}_{3}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms}\right),$ (51)
where “two other terms” means terms in which the role of ${\mathbf{p}}_{1}]$ is played by ${\mathbf{p}}_{2}]$ and ${\mathbf{p}}_{3}]$. One can again make contact with the order $(k_{\rm F}a_{0})^{3}$ contribution of the particle-hole diagram of Figure 2 to the ground state energy density $E/V$ computed in [14, 15] where it was given as a sum of two functions $K_{1}$ and $K_{2}$ (Eq. (17) in [14]), by using the algebraic identity
$N_{+-}^{{\mathbf{p}}_{1}}~\!{N_{-+}^{{\mathbf{p}}_{2}}-N_{+-}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{N_{-+}^{{\mathbf{p}}_{3}}-N_{+-}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms}\phantom{aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa}$ (52)
$=N_{+-}^{{\mathbf{p}}_{1}}~\!{N_{-+}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{N_{-+}^{{\mathbf{p}}_{3}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+N_{-+}^{{\mathbf{p}}_{1}}~\!{N_{+-}^{{\mathbf{p}}_{2}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}~\!{N_{+-}^{{\mathbf{p}}_{3}}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}]}+{\rm two~other~terms},$
[it is in fact the identity (35) but written with different symbols]. Using the symmetry of this expression allows to write the expression for $F^{(3)ph}/V$ as the sum of two terms that in the zero temperature limit reproduce the two terms, $K_{1}$ and $K_{2}$, which in Refs. [14, 15] represented the contribution of the second diagram of Fig. 2 to $E/V$.
Refer to caption Refer to caption
Figure 5: Dependence on the gas parameter $k_{\rm F}a_{0}$ of the resumed contributions of the particle-particle diagrams given by the expressions (47) (dashed blue lines), of the particle-hole diagrams given by (60) (dotted red lines) and of their sum (solid green lines) for zero temperature and two values of the polarization $P$.
In order to properly treat the singularities in (51) we again make the substitutions ${\mathbf{p}}_{l}]\rightarrow[{\mathbf{p}}_{l}]+il\epsilon$ which allow to profit from the symmetry of the three terms of (51). One then, similarly as in the case of $F^{(3)pp}$, obtains
${F^{(3)ph}\over V}=-C_{0}^{3}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!N_{+-}^{{\mathbf{p}}_{1}}\!\left[\left({\rm P}\!\int_{{\mathbf{p}}_{2}}{\{{\mathbf{p}}_{2}\}\over[{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}]}\right)^{2}+{1\over 3}\left(i\pi\!\!\int_{{\mathbf{p}}_{2}}\!\{{\mathbf{p}}_{2}\}~\!\delta([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}])\right)^{2}\right].$~ (53)
Making now the changes of the variables: first ${\mathbf{p}}_{1}+{\mathbf{q}}={\mathbf{k}}_{1}$, and ${\mathbf{p}}_{1}=-{\mathbf{k}}_{2}$, and then in the $n_{+}$ terms of $\{{\mathbf{p}}_{2}\}$ the change ${\mathbf{p}}_{2}+{\mathbf{k}}_{1}+{\mathbf{k}}_{2}=-{\mathbf{p}}$ and ${\mathbf{p}}_{2}={\mathbf{p}}$ in the $n_{-}$ terms of $\{{\mathbf{p}}_{2}\}$, and then taking explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the expression for $F^{(3)ph}/V$ can be written in the form
${F^{(3)ph}\over V}=-C_{0}^{\rm R}\!\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\![1-n_{+}({\mathbf{k}}_{1})]~\!n_{-}({\mathbf{k}}_{2})\left[K^{2}({\mathbf{k}}_{1},{\mathbf{k}}_{2})+{1\over 3}\left(iK_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})\right)^{2}\right],$ (54)
in which the dimensionless functions $K$ and $K_{\delta}$ are given by the integrals
[…]
$\left.+\int_{0}^{\infty}\!dp~\!p~\!n_{-}(p)\ln\!\left|{p-\Delta_{2}\over p+\Delta_{2}}\right|\right),$ (55)
$K_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})=\pi~\!{C_{0}^{\rm R}m_{f}\over(2\pi)^{2}\hbar^{2}|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}\left(\int_{|\Delta_{1}|}^{\infty}\!dp~\!p~\!n_{+}(p)-\int_{|\Delta_{2}|}^{\infty}\!dp~\!p~\!n_{-}(p)\right),$ (56)
in which
$\Delta_{1}\equiv{{\mathbf{k}}_{1}\!\cdot\!({\mathbf{k}}_{1}+{\mathbf{k}}_{2})\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!,\phantom{aaa}\Delta_{2}\equiv{{\mathbf{k}}_{2}\!\cdot\!({\mathbf{k}}_{1}+{\mathbf{k}}_{2})\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!.$ (57)
Again, the limits of the integrals in $K_{\delta}$ are determined by the condition that the zeros of the arguments of the Dirac deltas treated as functions of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lie between 1-1 and +1+1. And again the function $K_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can be written down in a closed analytical form - see below.
Refer to caption Refer to caption
Figure 6: As in Figure 5 but for $T=0.2~\!T_{\rm F}$.
The expression (54) is finite although this is not immediately obvious: while the integrals defining the functions $K$ and $K_{\delta}$ are clearly (ultraviolet) finite owing to the presence of the distribution functions $n_{+}(p)$ and $n_{-}(p)$, the integral over $d^{3}{\mathbf{k}}_{1}$ has no such exponentially suppressing factor. In addition, the presence of the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of the functions $K$ and $K_{\delta}$ seem to imply the potential presence of a singularity in the integral over the directions of ${\mathbf{k}}_{1}$ (for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$). To prove the finitness of (54) we will first analyze the behavior of the difference $K^{2}-K_{\delta}^{2}/3$ in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$. To this end it is helpful to split the functions $K$ and $K_{\delta}$ into $K_{+}+K_{-}$ and $K_{\delta+}+K_{\delta-}$ (the subscripts $\pm$ corresponds to the $n_{\pm}$ distribution functions). The hardest to see is the convergence of the integral over $d^{3}{\mathbf{k}}_{1}$ involving the factor $K_{-}^{2}-K^{2}_{\delta-}/3$. The dangerous term $n_{-}({\mathbf{k}}_{2})[K_{-}^{2}-K^{2}_{\delta-}/3]$ comes from the term
$\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\int_{{\mathbf{p}}_{2}}\!\int_{{\mathbf{p}}_{3}}\left\{{n_{-}({\mathbf{p}}_{1})~\!n_{-}({\mathbf{p}}_{2})~\!n_{-}({\mathbf{p}}_{3})\over([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{2}])([{\mathbf{p}}_{1}]-[{\mathbf{p}}_{3}])}+{\rm two~other~terms}\right\},$
in (51). This, however, is algebraically zero [just as the integrand of (36)] and, therefore, the factor $K_{-}^{2}-K^{2}_{\delta-}/3$ must be zero also. As any inaccuracy in the numerical evaluation of the integrals (55) and (56) could lead to a nonzero difference $K_{-}^{2}-K^{2}_{\delta-}/3$ and thus to a (fake) nonconvergence of the integration over $d^{3}{\mathbf{k}}_{1}$, in the term with unity arising from $1-n_{+}({\mathbf{k}}_{1})]$ we simply replace $K^{2}-K^{2}_{\delta}/3$ by $K_{+}^{2}+2K_{+}K_{-}-(K_{\delta+}^{2}+2K_{\delta+}K_{\delta-})/3$. The rest of the terms are separately integrable in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$:
$K_{\delta+}\propto{1\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!\ln\!\left(1+e^{-\beta(\hbar^{2}\Delta_{1}^{2}/2m-\tilde{\mu}^{(0)}_{+})}\right)\approx{1\over|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|}~\!e^{-\beta(\hbar^{2}\Delta_{1}^{2}/2m-\tilde{\mu}^{(0)}_{+})}~\!,$
because $\Delta^{2}_{1}$ grows like ${\mathbf{k}}_{1}^{2}$ as $|{\mathbf{k}}_{1}|\rightarrow\infty$. This secures the convergence of the integrals of the terms $K_{\delta+}^{2}$ and $2K_{\delta+}K_{\delta-}$. Similarly, it can be estimated that the integral over $p$ in $K_{+}$ behaves as $1/\Delta_{1}$ when $|{\mathbf{k}}_{1}|\rightarrow\infty$. Since each of the $K_{\pm}$ functions has the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of it, the term $(K_{+})^{2}$ behaves for $|{\mathbf{k}}_{1}|\rightarrow\infty$ as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\cdot{\mathbf{k}}_{2})^{2}$ and this secures the convergence of the integration over $d^{3}{\mathbf{k}}_{1}$. The term $2K_{+}K_{-}$, instead, behaves only as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\!\cdot\!{\mathbf{k}}_{2})|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, but the integration over the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$ kills the term of order $1/|{\mathbf{k}}_{1}|^{3}$ and the remaining terms are already integrable. However, this only power-like suppression of the integration of the term $K_{+}^{2}+2K_{+}K_{-}$ makes numerical evaluation of the particle-hole contribution to the free energy more difficult and therefore potentially less accurate than the evaluation of the particle-particle one. As to the potentially singular factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, one should first notice that the original expressions (48) and (49) as well as similar formulas giving the contributions of the $N$-th order particle-hole rings do not contain such singularities. They are due to the symmetrizations needed to arrive at the final formulas and must therefore cancel out [like the spurious singularities of the integrand in (36)] even if it is not directly evident. In the third order the finiteness of (54) can be seen as follows: since the integrals in the definitions of the $K_{\pm}$ and $K_{\delta\pm}$ functions are finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$, the singularities of (54) have essentially the form $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{2}^{2}+2|{\mathbf{k}}_{1}||{\mathbf{k}}_{2}|\xi)$ where $\xi$ is the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$. In the third order contribution considered here, after integration over $\xi$ they give rise to terms $\ln\!||{\mathbf{k}}_{1}|-|{\mathbf{k}}_{2}||$ which are integrable. [^5] In fact for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$ the factors $\Delta_{1}$ and $\Delta_{2}$ behave as $\sqrt{1+\xi}$ and it can be checked (numerically) that the integrals in the functions $K_{+}$ and $K_{-}$ vanish then as $\sqrt{1+\xi}$, so these functions have finite limits. The singularities reside only in the terms involving the functions $K_{\delta\pm}$. In the numerical evaluation of the integrals in (54) and (60) it is however sufficient to impose a cutoff $|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|>\kappa$ and check that the results stabilize as $\kappa$ approaches zero. Thus the expression (54) is finite and we have checked that evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) it reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-hole diagram of Figure 2 obtained in [14, 15].
Refer to caption Refer to caption
Figure 7: As in Figure 5 but for $T=0.5~\!T_{\rm F}$.
Using the same tricks as previously the contribution to the free energy of the infinite series of Feynman diagrams composed of $N$-fold rings of the particle-hole $B$-blocks of Figure 1 can be summed in a closed form. The order $(k_{\rm F}a_{0})^{N}$ term of this series is
${F^{(N)ph}\over V}=-(-1)^{N+1}~\!{C_{0}^{N}\over N}\!\int_{\mathbf{q}}\!\int_{{\mathbf{p}}_{1}}\!\dots\!\int_{{\mathbf{p}}_{N}}\sum_{n=1}^{N}N_{+-}^{{\mathbf{p}}_{n}}\left(\prod_{j\neq n}^{N}{\{{\mathbf{p}}_{j}\}\over[{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}]+i(n-j)\epsilon}\right),$ (58)
(apart from the extra minus originating from the minus sign in the identity $\{{\mathbf{p}}\}=-N^{\mathbf{p}}_{+-}(1-e^{\beta[{\mathbf{p}}]})$, the origin of the rest of the prefactor is the same as in the case of $F^{(N)pp}$) and, repeating the steps one arrives at the formal sum
$\sum_{N=1}^{\infty}{F^{(N)ph}\over V}=-~\!C_{0}^{\rm R}\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\![1-n_{+}({\mathbf{k}}_{1})]~\!n_{-}({\mathbf{k}}_{2})~\!{{\rm arctan}(K_{\delta}/(1-K))\over K_{\delta}}~\!.$ (59)
From this sum one has to subtract two first terms of the series: there is no order $C_{0}^{\rm R}$ particle-hole diagram and the order $(C_{0}^{\rm R})^{2}$ term is already (recall that the order $C_{0}^{2}$ contribution of the diagram shown in Fig. 1 can be written either as a convolution of two $A$-blocks or of two $B$-blocks shown in the same figure), in the properly renormalized form, included in (47). Thus the final form of the resummed contributions of the particle-hole ring diagrams is
${F^{ph}\over V}=-C_{0}^{\rm R}\int_{{\mathbf{k}}_{1}}\!\int_{{\mathbf{k}}_{2}}\![1-n_{+}({\mathbf{k}}_{1})]~\!n_{-}({\mathbf{k}}_{2})\left[{{\rm arctan}(K_{\delta}/(1-K))\over K_{\delta}}-1-K\right].$ (60)
One has to comment again on the finitness of the expression (60). The singularities related to the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ are now harmless because the function $K$ is, as remarked, finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$ and the singular function $K_{\delta}$ is now in the denominator and under the arctan function. As to the ultraviolet finiteness of (60), the integral of the factor $K$ explicitly subtracted in the square brackets is ultraviolet divergent being simply equivalent to the divergent expression (28). To ensure proper cancellation of this term in (60) for large $|{\mathbf{k}}_{1}|$ and/or $|{\mathbf{k}}_{2}|$ we expand the arctan function in $K_{\delta}$ and $K$ up to the sixth order (we have checked that taking more terms of the expansion does not change the result appreciably). The ultraviolet divergent terms linear in $K$ then disappear and, moreover, the expansion allows to implement the discussed trick with replacing $K^{2}-K_{\delta}^{2}/3$ by $K^{2}_{+}+2K_{+}K_{-}-(K^{2}_{\delta+}+2K_{\delta+}K_{\delta-})$.
In Ref. [15] we have compared the order $(k_{\rm F}a_{0})^{3}$ contributions to the ground state energy (i.e. to the free energy $F$ for zero temperature) of the particle-particle and of the particle-hole diagrams (shown in Fig. 2) in the entire range of the polarization $P$ and found that the second one is not much smaller than the first one. In Figs. 5-7 we compare the magnitudes of the resummed contributions (47) and (60) to the free energy as functions of $k_{\rm F}a_{0}$ for three different temperatures and two representative values of the system’s polarization $P$. We also plot the sum of these two contributions. It again follows that for $k_{\rm F}a_{0}\sim 1$ the resummed contribution of the particle-hole diagrams is not much smaller than that of the particle-particle ones and is of the opposite sign. It is also clear that in this most important region the sum of the two contributions is significantly distorted compared to the resummed contribution of the particle-particle diagrams. This raises the question of what impact the resummed contribution of the particle-hole ring diagrams has on the results reported in the Refs. [18, 19, 30].
Refer to caption Refer to caption
Figure 8: The difference $(F(P)-F(0))^{pp+ph}/N$ (in units $(3/5)\varepsilon_{\rm F}$ for $T=0$ and $T=0.2~\!T_{\rm F}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values (indicated in the panels) of the gas parameter $k_{\rm F}a_{0}$.
Figures 8 and 9, analogous to Figures 3 and 4 illustrate the consequences of adding the resummed contribution (60) of the particle-hole ring diagrams to the free energy. The result is dramatic: the phase transition to the ordered state discussed in Refs. [18, 19, 30] completely disappears! Thus, even if the selection of the particle-particle ring diagrams as giving the dominant contribution can be (partially) justified by invoking the arguments, given in Ref. [35], based on using $1/2^{D/2}$ where $D$ is the number of space dimensions, as the expansion parameter, they in practice do not turn out to be really dominant: the contribution of other subsets of diagrams (the number of such subsets beginning at a given order of the expansion grows with the order number) can, as our results show, change qualitatively the behavior of the thermodynamic potentials of the system of fermions close to the Feshbach resonance.
7. Conclusions
We have applied the systematic thermal (imaginary time) perturbative expansion to the effective (low energy) field theory to compute the free energy of the gas of interacting (nonrelativistic) spin $1/2$ fermions for arbitrary values of the gas polarization and temperatures not exceeding the Fermi temperature. We have shown how to circumvent the technical problem which previously prevented us from immediately extending such a computation beyond the second order in the gas parameter $k_{\rm F}a_{0}$ and have given explicit formulas for the order $(k_{\rm F}a_{0})^{3}$ contributions to the system’s free energy. It turned out that the analytical part of this computation is more transparent and easier than the corresponding direct computation of the ground state energy based on the formula (14) which gives only the zero temperature limit of the results obtained with the help of the thermal expansion. (Of course, numerical evaluation of the resulting expressions for a nonzero temperature is considerably more involved than for $T=0$).
Refer to caption Refer to caption
Figure 9: As in Figure 8 but for $T=0.3~\!T_{\rm F}$ and $T=0.5~\!T_{\rm F}$.
To obtain the complete order $(k_{\rm F}R)^{3}$ contribution to the free energy of the gas of fermions interacting through a truly repulsive spin-independent two-body potential (characterized by a length scale $R$) one would have to add the contributions arising from the operators of lower length dimension in the effective theory interaction term because such potentials naturally give the $p$-wave scattering length $a_{1}$ and the $s$-wave effective range $r_{0}$ comparable to the $s$-wave scattering length $a_{0}$. Instead of doing this, in this work we have profited from the simple structure of the contributions of the particle-particle and particle-hole ring diagrams and managed to give simple formulas for their contributions to the free energy resummed to all orders in the gas parameter $k_{\rm F}a_{0}$. The obtained formulas therefore apply rather to cold gases of fermionic atoms (interacting through attractive potentials) close to the Feshbach resonance where their $s$-wave scattering length $a_{0}$ is made positive and much larger than the remaining scattering lengths and effective radii. Using these formulas we have first checked that including only the contributions of the particle-particle rings we reproduce for zero temperatures all the results obtained in Ref. [18]. In particular we confirm that in this approximation at zero temperature the transition to the ordered phase occurs for $k_{\rm F}a_{0}=0.79$ and that it is continuous. These results seem to agree well with the results of the dedicated Monte Carlo computations. Our formula would allow to study also the thermal profile of the transition, if it really existed.
However we have found that the phase transition to the ordered state completely disappears after including into the free energy the resummed contribution of the particle-hole ring diagrams - the minimum of the free energy is always for zero polarization. This may at least indicate that the agreement of the critical value of the gas parameter $k_{\rm F}a_{0}$ found in Refs. [18, 19, 30] with its value obtained from the Monte Carlo simulations (done with attractive potentials tuned so that $a_{0}$ is positive and large) is just accidental. It may however also indicate that there is indeed no transition (in the metastable state) to the magnetically ordered phase if the true interaction is attractive (but $a_{0}$ is made positive and large by approaching a Feshbach resonance). The vanishing at zero temperature of the inverse spin susceptibility ($1/\chi\propto(\partial^{2}F/\partial P^{2})_{P=0}$) for some value of the gas parameter found in the Monte Carlo simulations (which in the case of attractive potentials are not as clean as in the repulsive case because of the need to exclude - in the finite volume and a finite number of fermions, that is without taking the thermodynamic limit - an overlap with the true ground state of the considered system of atoms) [21] from which one infers its existence and its continuous character would then be misleading (the second order approximation to $F$ predicts analytically vanishing of the inverse spin susceptibility at $k_{\rm F}a_{0}=1.058$ whereas the transition is first order and occurs for $k_{\rm F}a_{0}=1.054$ [18]). If this hypothesis is true it would provide yet another (in addition to the formation of atomic dimers) reason for the failure to simulate the itinerant ferromagnetism with the help of cold atoms. In an case it is clear that more theoretical studies are needed to clarify the situation. An obvious step in this direction would be extending the computation of ref. [10] of the order $(k_{\rm F}a_{0})^{4}$ contribution to nonequal densities of spin up and spin down fermions.
Data Availability Statement
The source code for this project is publicly available on Zenodo [36].
Acknowledgements
This research was funded, in whole or in part, by l’Agence Nationale de la Recherche (ANR), project ANR-23-CE31-0019..